### v2

# Locality in Residuated Lattice Structures

Abstract.
Many-valued models generalise the structures from classical model theory by defining truth values for a model with an arbitrary algebra. Just as algebraic varieties provide semantics for many non-classical propositional logics, models defined over algebras in a variety provide the semantics for the corresponding non-classical predicate logics. In particular models defined over varieties of residuated lattices represent the model theory for first-order substructrual logics.
In this paper we study the extent to which the classical locality theorems from Hanf and Gaifman hold true in the residuated lattice setting. We demonstrate that the answer is sensitive both to how locality is understood in the generalised context and the behaviour of the truth-defining algebra. In the case of Hanf’s theorem, we will show that the theorem fails for the natural understanding of local neighbourhoods, but is recoverable in one special case for well-connected residuated lattices. For Gaifman’s theorem, rather than consider Gaifman normal forms directly we focus on the main lemma of the theorem from textbook proofs. We prove that for a number of different understandings of locality, provided the algebra is well-behaved enough to express locality in its syntax, this main lemma can be recovered. In each case we will see that importance of an order-interpreting connective which creates a link between the modelling relation between models and formulas and the valuation function from formulas into the algebra.
Key words and phrases:
optional comma separated list of keywords
thanks, optional.
1. Introduction
An important limitation of classical first-order is its locality. Whether a first-order formula is modelled in a given structure depends only on the behaviour of the model at a ‘local’ level, with the consequence that first-order logic cannot express ‘global’ properties. This is at its most transparent in graphs where the natural notions of distance between points makes precise the idea of local, and locality results provide an effective method to show some natural graph concepts such as connectivity or acyclicity are inexpressible in logical terms [Lib04, Section 3.6]. In arbitrary signatures locality is made precise through the Gaifman graph of a model, allowing the natural notion of distance in a graph to define one for any structure [EF95, Section 1.4]. Once distance is understood a multitude of locality results can be expressed, the two most significant being Hanf locality [Han65] and Gaifman locality [Gai82]. Hanf locality links the equivalence of sentences up to a given quantifier rank to the number of isomorphic local substructures; Gaifman locality states that any first-order formula is equivalent to one restricted to ‘local quantification’. Locality results are especially interesting in the context of finite model theory due to the topic of database queries for which locality results provide an effective method of proving a given query is not first-order definable[Lib04].
The utility of locality results in classical finite model theory motivates understanding how similar theory could be developed in other model theoretic contexts. Previous work in this vein comes from Bizière, Grädel and Naaf [BGN23] who investigated locality for finite models defined over ordered semirings, and from Wild, Schröder, Pattinson and König [WSPK18] who remarked on locality for the first-order fragment corresponding to a fuzzy modal logic. Here we aim to do the same for finite models defined over residuated lattices. Varieties of residuated lattices provide the algebraic semantics for a multitude of many-valued logics encompassing many of the most well-studied substructural logics [GJKO07]. Specific varieties have been the setting of much work regarding non-classical model theory, including the development of back-and-forth criteria [DGCN18], the status of amalgamation theorems [BCDN19] or 0-1 laws [BCN25]. Notably, neither the semiring and residuated lattice setting subsumes the other. Semiring semantics takes standard classical syntax and generalises the behaviour of the conjunction and disjunction connective but lacks a primitive notion of implication or negation. The result is that our investigation develops quite differently from that of Bizière et al.
Our choice to work with residuated lattices is far from arbitrary, rather it reflects a common pattern of research in non-classical logic reflected in our investigation. Namely, we identity a topic from classical model theory, adjust the core concepts for the generalised setting and attempt to recover a proof in as weak a setting as possible. As problems arise the behaviour of models is strengthened, adopting the main features of classical logic used in the original proof. Ideally once a theorem is recovered those features are shown to be necessary. In this sense to study non-classical model theory is also to study classical model theory identifying the exact critical features of classical models that underlie a particular theorem. When it comes to locality, an ongoing theme will the importance of the presence of an order-defining connective, that is an algebraic term $t(x,y)$ that takes a ‘true’ value exactly when $x$ is less than or equal to $y$ in the lattice order. In residuated lattices this is a critical property of the left and right residuums of the strong conjunction, or in the commutative case the implication connective $\rightarrow$. The presence of this connective creates a link between the usual modelling relation $\models$ and the valuation function ||||||-||| | - | | which constitute the two central relationships between syntax and semantics, and from this link the more substantive syntax-semantic relationships of isomorphism types and locality theorems are then recovered.
Exploring these questions in the residuated lattice setting the initial question is how to give generalisations of the concepts needed to express locality. As is not uncommon in many-valued model theory, there are multiple sensible generalisations united in some basic properties and relationships. In the case of locality, what emerges is that those basic properties and relationships are sufficient to recover much of the theory, with the proof work itself following through as it does classically. This determines the shape of our investigation. In Section 2 we properly introduce many-valued models as well as the algebras they are defined relative to. We then briefly detail some of the critical model theory for these models that will be necessary for the main investigation, including highlighting the particularly important subclasses of standard expansions and witnessed models. Our main investigation is split into two sections, Section 3 focusing on Hanf locality and Section 4 on Gaifman locality. In each case we establish a variant of the locality results for models defined over a residuated lattice with some additional constraints. We then compare and contrast our work with residuated lattices with the investigation for semirings in Section 5. Finally, in Section 6 we briefly see how our locality results can be used in similar manner to the classical results to provide a method of showing queries for the considered models are not definable, before concluding with some remarks about further study.
As in the classical setting, locality is investigated only for relational languages, possibly with object constants [EF95, Lib04]. Furthermore, as is common in the many-valued setting, we restrict to models defined over a fixed algebra (see also [BCDN19, DV19]).
2. Preliminaries
We start with a brief introduction our objects of study - many-valued models. Although we will focus solely on finite models in our main investigation we introduce our models in a general manner. Many-valued models consist of two components, a non-empty set equipped with interpretations of non-logical symbols akin to familiar classical models, and an algebra that is used to define truth. Naturally, the properties of the structures is tied to the choice of algebras they are defined over. Different motivations lead one to be concerned with different classes of models. The form of our investigation is an attempt to understand for which truth-defining algebras some interesting model theory can be recovered - here the theory in question is locality. We choose to work at a level of generality that can be applied cleanly to a wide range of more specific motivated settings whilst also making transparent the essential properties the locality results are sensitive to.
{defi}
A residuated lattice is an algebra $A$ in signature $\mathcal{L}=\langle\land,\lor,\cdot,\mathord{\reflectbox{/}},/,1\rangle$ such that:
  • $\langle A,\land,\lor\rangle$ is a lattice.
  • $\langle A,\cdot,1\rangle$ is a monoid
  • The residuation properties hold for $\langle A,\cdot,\mathord{\reflectbox{/}},/\rangle$; for all $a,b,c\in A$
    $a\leq c/b\text{ iff }a\cdot b\leq c\text{ iff }b\leq a\mathord{\reflectbox{/}}c.$
A residuated lattice $A$ is well-connected iff for all $a,b\in A$ $a\lor b\geq 1$ iff $a\geq 1$ or $b\geq 1$.
Residuated lattices provide an interesting setting for many-valued model theory as they provide the algebraic semantics for substructural logics, a well established class of propositional logics containing most of the well-studied particular logics in the literature, including Boolean algebras, Heyting algebras, MV-algebras and lattice ordered groups [GJKO07, CHN11, Section 3.4]. The models defined over these algebras provide the semantics to the first-order predicate logics extending their respective propositional logic [CN21, Chapter 7]. The consideration of well-connected lattices will be essential to have the appropriate behaviour of the existential quantifier and is partially justified due to all (finitely) subdirectly irreducible ((F)SI) residuated lattices being well-connected [GJKO07, Lemma 5.29]. This gives the logical viewpoint of our investigation, to what extent does the locality of classical logic extend to substructural logics.
We can immediately highlight the basic properties of well-connected residuated lattices that motivate their use for the algebras in our models. Firstly is that the residual operators interpret the algebraic order,
$1\leq b/a\text{ iff }a\leq b\text{ iff }1\leq a\mathord{\reflectbox{/}}b.$
Secondly is that our lattice connectives are stable around $\geq 1$,
[…]
Let us now introduce our basic syntax and semantics. {defi} A predicate language $\mathcal{P}$ is a pair $\langle R,R\rangle$ where $F$ is a non-empty set of function symbols and $R$ a set of relation symbols ($R$, $F$ are assumed disjoint) all of which come with an assigned natural number - its arity. The relation symbols of arity zero are called truth constants and function symbols of arity zero are called object constants.
Let $\mathcal{L}$ be a propositional/algebraic signature $\mathcal{L}$ and $Var$ a countable set of variables. We define the sets of $\mathcal{L},\mathcal{P}$-terms, $\mathcal{L},\mathcal{P}$-atomic formulas $\mathcal{L},\mathcal{P}$-formulas inductively just as in classical logic except we naturally insist that the formulas are closed under all connectives from $\mathcal{L}$. We denote the set of $\mathcal{P}$-formulas by $Fm(\mathcal{P})$. We similarly define notions such as free occurrence of a variable, open formula, substitutability and sentence as in classical model theory.
We always assume that we have the equality symbol $=$ in our predicate language.
Remark 1.
Whilst we present our basic definitions for signatures including function symbols, our investigation is restricted to relational languages, i.e. those without function symbols. Additionally, as the algebraic signature is usually clear from context we frequently omit reference to it and talk of $\mathcal{P}$-terms, formulas, sentences etc.
{defi}
Let $\mathcal{P}$ be a predicate language. A $\mathcal{P}$-structure is a pair $\langle A,\mathcal{M}\rangle$ where $A$ is a non-trivial (well-connected) residuated lattice and $\mathcal{M}$ is a triple $\langle M,\langle F^{M}\rangle_{F\in\mathcal{P}}\langle R^{M}\rangle_{R\in\mathcal{P}}\rangle$ where:
  • $M$ is a non-empty set.
  • $F^{M}\colon M^{n}\rightarrow M$ for each $n$-ary function symbol $F\in\mathcal{P}$ ($F^{M}\in M$ when $n=0$).
  • $R^{M}\colon M^{n}\rightarrow M$ for each $n$-ary relation symbol $R\in\mathcal{P}$ ($R^{M}\in A$ when $n=0$).
The set $M$ is called the domain of $\mathcal{M}$ and the mappings $F^{M},R^{M}$ the interpretations of function and predicate symbols in $\mathcal{M}$.
Remark 2.
We frequently drop reference to $A$ and speak simply of $\mathcal{P}$-models $M$ either when $A$ is clear from context or when we are restricting attention to models defined over the same fixed algebra. In this case we frequently use $M$ interchangeably for the model and its underlying domain.
In the definition of structure we see an important difference between propositional truth-constant $d\in\mathcal{L}$ and an interpreted nullary relation symbol $R\in\mathcal{P}$. The former will be interpreted by the same truth value $d\in A$ for each structure $(A,M)$ whereas the latter is interpreted by some truth value $P^{M}\in M$ in each structure separately.
The insistence on $A$ being non-trivial is needed to have a coherent interpretation of the equality symbol.
{defi}
Let $\mathcal{M}$ be a $\mathcal{P}$-structure. We define an $M$-valuation as a mapping $v\colon Var\rightarrow M$. Given an $M$-valuation $v$, an object variable $x$ and an element $m\in M$ we denote by $v_{x=m}$ the $M$-valuation defined as:
$v_{x=m}:=\begin{cases}m&\text{if }y=x\\ v(y)&\text{otherwise}\end{cases}$
We extend an $M$-valuation $v$ to all $\mathcal{P}$-terms by setting for each $n$-ary $F\in\mathcal{P}$:
$v(F(t_{1},...,t_{n})).$
We define the valuation function $||-||\colon Fm(\mathcal{P})\rightarrow A$ of the $\mathcal{P}$-formulas for a given $M$-valuation $v$ recursively as follows:
$||t_{1}=t_{2}||^{M}_{v}$ =$\begin{cases}1&\text{if }v(t_{1})=v(t_{2})\\ a&\text{if }v(t_{1})\not=v(t_{2}),a\not=1\end{cases}$
$||P(t_{1},..,t_{n})||^{M}_{v}$ $=R^{M}(v(t_{1}),...,v(t_{n})),\text{ for each }n\text{-ary }P\in\mathcal{P},$
$||\circ(\varphi_{1},...,\varphi_{n})||^{M}_{v}$ $=\circ^{A}(||\varphi_{1}||^{M}_{v},...,||\varphi_{n}||^{M}_{v}),\text{ for each }n\text{-ary }\circ\in\mathcal{L},$
$||\forall x\varphi(x)||^{M}_{v}$ $=inf_{\leq}\{||\varphi(x)||^{M}_{v_{x=m}}:m\in M\},$
$||\exists x\varphi||^{M}_{v}$ $=sup_{\leq}\{||\varphi(x)||^{M}_{v_{x=m}}:m\in M\}.$
If the infimum or supremum does not exist we take the truth-value of such a formula to be undefined. We say that a $\mathcal{P}$-structure is safe iff $||\varphi||^{M}$ is defined for every formula and reserve the term $\mathcal{P}$-model for safe $\mathcal{P}$-structures.
Remark 3.
The use of $a\in A\setminus\{1\}$ to define equality is only important insofar as it means equality is crisp. In more specific circumstances there are usually a natural value for inequality to take, e.g. when $A$ is bounded $a$ would be taken as the lower bound $\bot$.
As often in the case of model theory, we usually suppress the direct consideration of the valuation $v$ and write as if the function $||-||^{M}$ is acting directly on $\mathcal{P}\cup M$ sentences, e.g. writing $||R(m_{1},...,m_{n})||^{M}$ in place of $||R(t_{1},...,t_{n})||^{M}_{v}$ where $||t_{i}||^{M}_{v}=m_{i}$.
We are exclusively interested in safe $\mathcal{P}$-structures, hence the reservation of the term models; indeed any finite $\mathcal{P}$-structure defined over a lattice is safe really eliminating the need for the distinction for finite structures. The distinction is mostly relevant when defining (infinite) $\mathcal{P}$-structures where we need to be careful to ensure they are safe, frequently this is done by choosing $A$ to be a complete lattice as then again any $\mathcal{P}$-structure defined over $A$ would be safe.
[…]
Defining the quantifier depth ($qd(-)$) of a $\mathcal{P}$-formula exactly as in classical logic we say that $M,N$ are $k$-strongly equivalent, denoted $M\equiv^{s}_{k}$ iff for all $\varphi\in Fm:qd(\varphi)\leq k$ $||\varphi||^{m}=||\varphi||^{N}$.
{defi}
Let $M,N$ be two $\mathcal{P}$-models defined over the same algebra $A$. A map $f\colon M\rightarrow N$ is a strong homomorphism iff:
for every $\text{for every }F\in\mathcal{P}\text{ and }\bar{m}\in M\ f[F^{M}(\bar{m})]=F^{N}(f(\bar{m})).$
[…]
Embeddings are injective strong homomorphism and isomorphisms are bijective strong homomorphisms. We say that $N$ is a substructure of $M$ iff there is an embedding $\iota\colon N\hookrightarrow M$.
We also introduce an essential tool for model theory - back-&-forth systems. These have been introduced for the many-valued setting in [DGCN18] for models defined over UL-algebras. In that discussion it is noted that the proofs can be easily generalised to apply to models over a wider choice of algebras and this includes residuated lattices. We make a slight adjustment to the definitions and results to simplify them given we will be working with a fixed defining algebra but the proof of the theorem is identical to [DGCN18, Theorem 22].
{defi}
Let $M,N$ be $\mathcal{P}$-models. We say that a partial mapping $r\colon M\rightarrow N$ is a partial isomorphism (p.iso) from $M$ to $N$ iff $r$ is an embedding on its domain, i.e. $r$ is injective and for every $R\in\mathcal{P}$ and $\bar{m}\in M:\bar{m}\in dom(r)$ $R^{M}(\bar{m})=R^{N}(r(\bar{m}))$.
We say $M$ and $N$ are finitely isomorphic, denoted $M\cong_{f}N$ iff there is a sequence $\langle I_{n}:n\in\omega\rangle$ with the following properties:
  1. (1)
    Every $I_{n}$ is a non-empty set of partial isomorphisms from $M$ to $N$ and $I_{n+1}\subseteq I_{n}$.
  2. (2)
    Forth: for every $r\in I_{n+1}$ and $m\in M$ $\exists r^{\prime}\in I_{n}:r\subseteq r^{\prime}$ and $m\in dom(r^{\prime})$.
  3. (3)
    Back: for every $r\in I_{n+1}$ and $n\in N$ $\exists r^{\prime}\in I_{n}:r\subseteq r^{\prime}$ and $n\in im(r^{\prime})$.
We say that $M$ and $N$ are $k$-finitely isomorphic, denoted $M\cong_{k}N$ iff there is such a sequence $I_{j}:j\leq k$
Theorem 4.
Let $\mathcal{P}$ be a finite relational language, $M,N$ be $\mathcal{P}$-models and $k\in\omega$. Suppose $M\cong_{k}N$. Then $M\equiv^{s}_{k}N$.
One can investigate these models as independent mathematical objects, working solely with the valuation function. [^1] Examples of investigations with this focus are [BGN23] and [HMV17]. However, the logical viewpoint motivates considering a $\models$ relation linking syntax and semantics in a less fine grained way.
{defi}
Let $A$ be a (well-connected) residuated lattice. We say that a $\mathcal{P}$-model $\mathcal{M}=(A,M)$ is a model of a $\mathcal{P}$-sentence $\varphi$ iff $||\varphi||^{M}\geq 1^{A}$. We also say that $\varphi$ is valid in $M$.
Given a subset of residuated lattices $\mathbb{K}$ the $\mathbb{K}$ based predicate logic in language $\mathcal{P}$ is the relation $\models^{\mathcal{P}}_{\mathbb{K}}$ defined for $\Gamma\cup\{\varphi\}\subseteq Fm(\mathcal{P})$ by:
$\Gamma\models^{\mathcal{P}}_{\mathbb{K}}\varphi\text{ iff for each }A\in\mathbb{K}\text{ and each }\mathcal{P}\text{-model }(A,M),(A,M)\models\Gamma\text{ implies }(A,M)\models\varphi.$
[…]
Notably, in the case of finite models these extend to the behaviour of $\forall$ and $\exists$ giving a limited form of witnessing:
[…]
This points towards a more ‘axiomatic’ version of our investigation; for any class of models defined over an algebra $A$ in signature $\mathcal{L}$ we could proceed on the assumption that there are some algebraic terms $t_{\land}(x,y)$ $t_{\lor}(x,y)$ and $t_{\leq}(x,y)$ with this behaviour relative to a given $\models$ relation. We will however proceed in a more concrete setting.
There are two subclasses of models that are important for our investigation. The first are the witnessed models. {defi} We say that a model $(A,M)$ is $\exists$-witnessed iff for all formulas $\exists x\varphi(x)$ there is an element $m\in M$ for which $\varphi(m)$ achieves the value of the supremum, i.e. $||\exists x\varphi(x)||^{M}=||\varphi(m)||^{M}$. We define $\forall$-witnessing similarly and say that $M$ is witnessed when it is both $\exists$ and $\forall$-witnessed.
Remark 5.
In the case of finite models $\exists$ and $\forall$ essentially become shorthand for strings of $\land$ and $\lor$ and thus whether a model is witnessed is determined by their behaviour. In particular, finite models defined over linear algebras are always witnessed, and as a partial converse if every $\mathcal{P}$-model over an algebra $A$ is witnessed then $A$ is linear.
The second are the standard expansions of models where we add fresh relation symbols to our language and whose intended interpretation are the truth values of the algebra. This lets us introduce syntactic truth constants at the predicate language level rather than in the algebraic signature. {defi} Let $A$ be a well-connected residuated lattice and $\mathcal{P}$ a predicate language. We expand the language $\mathcal{P}$ to one containing a 00-ary relation symbol for each element $a\in A$, i.e. $\mathcal{P}^{\ast}:=\mathcal{P}\cup A$. For each $\mathcal{P}$-model $M$ defined over $A$ its standard expansion $M^{\ast}$ is the $\mathcal{P}^{\ast}$-model for which $||a||^{M}=a$ for all $a\in A$.
Remark 6.
A common proof strategy in many-valued model theory is to work with the standard expansions of models before pulling the results back to the $\mathcal{P}$-reducts. This naturally requires a bridge result between the behaviour of a model and its standard expansion. The first and most basic is simply that for any $\mathcal{P}$-formula $\varphi(\bar{x})$ and $\bar{m}\in M$ $||\varphi(\bar{m})||^{M}=||\varphi(\bar{m})||^{M^{\ast}}$. For this reason we usually suppress reference to the standard expansion in the ||||||-||| | - | | notation. The second we use frequently is that $M\cong_{k}N$ iff $M^{\ast}\cong_{k}N^{\ast}$ which follows simply by the observation that any $\mathcal{P}$ back-&-forth system for $M$ and $N$ is immediately also a $\mathcal{P}^{\ast}$ system from $M^{\ast}$ and $N^{\ast}$ and vice versa.
3. Hanf Locality
We now turn to our first locality result - Hanf locality. Hanf’s theorem for classical logic is concerned with isomorphism types, a kind of syntactic encoding of b&f-systems, applied to local neighbourhoods. Our first task then is to set up a coherent many-valued counterpart to both these concepts. Throughout this section we define our models over a fixed a well-connected residuated lattice $A$.
We start with the isomorphism types. An extension of the notion of diagrams, at a base level the idea is to define a formula that encodes partial isomorphisms and then extend these with formulas that encode the back and forth conditions. For our models, by making use of the standard expansions we can make an adjustment to the usual classical encoding as described in [EF95].
{defi}
Let $\mathcal{P}$ be a finite relational language and $A$ be a well-connected residuated lattice. Let $v_{1},...,v_{s}$ be variables. For each finite $\mathcal{P}$-model $M$, $\bar{m}\in M$ and $k\in\omega$ we define the $\mathcal{P}^{\ast}$-formula $\varphi^{k}_{M,\bar{m}}(\bar{v})$ inductively as follows:
$\varphi^{0}_{M,\bar{m}}(\bar{v})$ :=$:=\bigwedge\{[R(\bar{v})\mathord{\reflectbox{/}}a]\land[R(\bar{v})/a]\in Fm:R(\bar{v})\in\mathcal{P}\text{ with arity }t\leq s\text{ and }||R(\bar{m})||^{M}=a\}.$
$\varphi^{k}_{M,\bar{m}}(\bar{v})$ $:=\bigwedge\limits_{m\in M}\exists v_{s+1}\varphi^{k-1}_{\bar{m}m}(\bar{v},v_{s+1})\land\forall v_{s+1}\bigvee\limits_{m\in M}\varphi^{k-1}_{\bar{m}m}(\bar{v},v_{s+1}).$
We call $\varphi^{k}_{M,\bar{m}}(\bar{v})$ the $k$-isomorphism type of $\bar{m}$ in $M$. We frequently drop reference to $M$ when it is clear from context.  [^3] Notably the definition of 00-isomorphism types is only sensitive to the behaviour of $\bar{m}$ in $M$ and independent from the rest of the model.
Remark 7.
The set used in $\varphi^{0}_{\bar{m}}$ is finite for any selection of $\bar{m}$ and moreover as $M$ is finite by a simple induction on $k$ for $s,k\geq 0$ the set $\{\varphi^{k}_{M,\bar{m}}:\bar{m}\in M^{s}$} is finite. Thus the formulas are all we defined.
Recall we have crisp equality in our language and so the 00-isomorphism types contain in particular $(v_{i}=v_{j})\mathord{\reflectbox{/}}1\land(v_{i}=v_{j})/1$ iff $m_{i}=m_{j}$ and $v_{i}=v_{j}\mathord{\reflectbox{/}}a\land v_{i}=v_{j}/a$ iff $m_{i}\not=m_{j}$.
We can quickly establish some basic facts about isomorphism types.
Lemma 8.
For any $\mathcal{P}$-model and $k\in\omega$ we have the following:
  1. a)
    $qd(\varphi^{k}_{\bar{m}}(\bar{v}))=k$,
  2. b)
    $M^{\ast}\models\varphi^{k}_{\bar{m}}(\bar{m})$,
  3. c)
    For any $N$ and $\bar{n}\in N$ $N^{\ast}\models\varphi^{0}_{\bar{m}}(\bar{n})$ iff $\bar{m}\mapsto\bar{n}$ is a partial isomorphism (p.iso).
Proof 3.1.
$a)$ is immediate from the syntax construction and $b)$ follows by an easy induction. For c) suppose if $N^{\ast}\models\varphi^{0}_{\bar{m}}(\bar{n})$. Firstly, the map $\bar{m}\mapsto\bar{n}$ is injective by our previous remark:
$m_{i}\not=m_{j}\Rightarrow v_{i}=v_{j}\mathord{\reflectbox{/}}\bot\land v_{i}=v_{j}/\bot\in\varphi^{0}_{\bar{m}}(\bar{v})\Rightarrow||n_{i}=n_{j}||^{N^{\ast}}=\bot\Rightarrow n_{i}\not=n_{j}.$
The only function symbols allowed are object constants $c\in\mathcal{P}$ and:
$c^{M}=m_{i}\in\bar{m}\Rightarrow c=v_{i}\mathord{\reflectbox{/}}\top\land c=v_{i}/\top\in\varphi^{0}_{\bar{m}}(\bar{v})\Rightarrow||c=n_{i}||^{N}=\top\Rightarrow c^{N}=n_{i}.$
Finally, letting $R\in\mathcal{P}$ and $\bar{m^{\prime}}\subseteq\bar{m}$ be such that $R^{M}(\bar{m^{\prime}})=m$ then $R(\bar{v^{\prime}})\mathord{\reflectbox{/}}a\land R(\bar{v^{\prime}})/a\in\varphi^{0}_{\bar{m}}(\bar{v})$ which implies $N^{\ast}\models\varphi^{0}_{\bar{m}}(\bar{n})\Rightarrow||R(\bar{n^{\prime}})\mathord{\reflectbox{/}}a\land R(\bar{n^{\prime}})/a||\geq 1$ and so $|R(\bar{n^{\prime}})||^{N}=a=||R(\bar{m^{\prime}})||^{M}$. Therefore $\bar{m}\mapsto\bar{n}$ is a p.iso.
For the converse, if $\bar{m}\mapsto\bar{n}$ is a p.iso then letting $R(\bar{v})$ be an atomic $\mathcal{P}$-formula with arity $t\leq s$ and $||R(\bar{m})||^{M}=a$ we have $|R(\bar{n}))||^{N}=||R(\bar{m})||^{M}=a$ and in particular $||R(\bar{n})\mathord{\reflectbox{/}}a\land R(\bar{n})/a||^{N}\geq 1$, i.e. $N^{\ast}\models\varphi^{0}_{\bar{m}}(\bar{n})$.
Remark 9.
An immediate consequence of $iii.$ is given any two tuples $\bar{m}\in M$ and $\bar{n}\in N$ they realise the same type iff $\bar{m}\mapsto\bar{n}$ is a partial isomorphism.
These help prove their defining behaviour - their encoding of b&f-systems.
Lemma 10.
Let $\mathcal{P}$ be a finite relational language and $A$ a well-connected residuated lattice. Let $M,N$ be finite models defined over $A$. Let $\bar{m}\in M$, $\bar{n}\in N$ and $k\in\omega$. The following are equivalent:
  1. i.
    $M,\bar{m}\cong_{k}N,\bar{n}$ with $\bar{m}\mapsto\bar{n}\in I_{k}$.
  2. ii.
    $N^{\ast}\models\varphi^{k}_{M,\bar{m}}(\bar{n})$.
  3. iii.
    $M,\bar{m}\equiv^{s}_{k}N,\bar{n}$.
Proof 3.2.
That $iii.$ implies $ii.$ is immediate from our previous lemma and noting that strong equivalence is preserved when moving to the standard expansion and Theorem 4 gives that $i.$ implies $iii.$. So suppose $N^{\ast}\models\varphi^{k}_{\bar{m}}(\bar{n})$, we describe how to construct the b-&-f system $\langle I_{j}\rangle_{j\leq k}$.
We define $I_{k}:=\{\bar{m}\mapsto\bar{n}\}$ and check that this is a p.iso. We proceed by induction on $k$ where $k=0$ is the first direction of part $c)$ of the previous lemma. Letting $k>0$ $\bar{m}\in M$ and $\bar{n}\in N$ with $N^{\ast}\models\varphi^{k}_{\bar{m}}(\bar{n})$ then in particular, taking an arbitrary $m\in M$ we have $N^{\ast}\models\exists v_{s+1}\varphi^{k-1}_{\bar{m}m}(\bar{n}v_{s+1})$ and by $A$ well-connected $\exists n\in N:N^{\ast}\models\varphi^{k-1}_{\bar{m}m}(\bar{n},n)$. Then by induction hypothesis $\bar{m}m\mapsto\bar{n}n$ is a p.iso and in particular so is $\bar{m}\mapsto\bar{n}$.
Now, for each $m\in M$ $N^{\ast}\models\exists v_{s+1}\varphi^{k-1}_{\bar{m}m}(\bar{n},v_{s+1})$ and so by $A$ well-connected $\exists n\in N:N^{\ast}\models\varphi^{k-1}_{\bar{m}m}(\bar{n},n)$. Thus, again by the previous lemma $\bar{m}m\mapsto\bar{n}n$ is a p.iso and so we may define the non-empty set of p.isos $I^{l}_{k-1}:=\{\bar{m}m\mapsto\bar{n}:N^{-}\models\varphi^{k-1}_{\bar{m}m}(\bar{n},n)\}$. Similarly, letting $n\in N$ we have in particular that $N^{\ast}\models\bigvee\limits_{m\in M}\varphi^{k-1}_{\bar{m}m}(\bar{n},n)$ and as $A$ is well-connected there is an $m\in M:N^{\ast}\models\varphi^{k-1}_{\bar{m}m}(\bar{n},n)$. Therefore we may once again define a non-empty set of p.isos $I^{r}_{k-1}:=\{\bar{m}m\mapsto\bar{n}n:N^{-}\models\varphi^{k-1}_{\bar{m}m}(\bar{n}n)\}$. We then define $I_{k-1}=I^{l}_{k-1}\cup I^{r}_{k-1}$.
We may repeat this process on each element of $I_{j}$ for $j<k$ and take the total union to define $I_{j-1}$ and in this way define our sequence $\langle I_{j}\rangle_{j\leq k}$. By construction each set is a non-empty set of p.isos and the sequence satisfies the b-&-f conditions, with $\bar{m}\mapsto\bar{n}\in I_{k}$ as required.
Later we will wish to apply this result directly to standard expansions of models. Whilst the expanded language $\mathcal{P}^{\ast}$ is not finite we can recover the equivalence with a little extra work.
Corollary 11.
Let $\mathcal{P}$ be a finite relational language and $A$ a well-connected residuated lattice. Let $M,N$ be finite models defined over $A$. Let $\bar{m}\in M,\bar{n}\in N$ and $k\in\omega$. The following are equivalent:
  1. i.
    $M,\bar{m}\cong_{k}N,\bar{n}$ with $\bar{m}\mapsto\bar{n}\in I_{k}$.
  2. ii.
    $M^{\ast},\bar{m}\cong_{k}N^{\ast},\bar{n}$ with $\bar{m}\mapsto\bar{n}\in I_{k}$.
  3. iii.
    $N^{\ast}\models\varphi^{k}_{M,\bar{m}}(\bar{n})$.
  4. iv.
    $M,\bar{m}\equiv^{s}_{k}N,\bar{n}$.
  5. v.
    $M^{\ast},\bar{m}\equiv^{s}_{k}N^{\ast},\bar{n}$.
Proof 3.3.
Our observation that any $\mathcal{P}$ back-&-forth system for $M$ and $N$ is also a $\mathcal{P}^{\ast}$ system for $M^{\ast}$ and $N^{\ast}$ combined with the previous lemma gives the equivalence of $i.$ through $iv.$ and by taking the $\mathcal{P}$ reduct $v.$ implies $iv.$ To conclude we show that $ii.$ implies $v.$. Let $\varphi\in Fm(\mathcal{P}^{\ast})$ be of quantifier depth at most $k$ and $\bar{a}$ be the finite list of the newly introduced nullary relation symbols occurring in $\varphi$. We consider the finite relational language $\mathcal{P}^{\prime}=\mathcal{P}\cup\{\bar{a}\}$ and the $\mathcal{P}^{\prime}$-reducts $M^{\prime}$ and $N^{\prime}$ of $M^{\ast}$ and $N^{\ast}$ respectively. Then $ii.$ implies that $M^{\prime},\bar{m}\cong_{k}N^{\prime},\bar{n}$ and we may apply our previous lemma to $M^{\prime},\bar{m}$ and $N^{\prime},\bar{n}$ to conclude $M^{\prime},\bar{m}\equiv^{s}_{k}N^{\prime},\bar{n}$. In particular $||\varphi||^{M^{\prime}}=||\varphi||^{N^{\prime}}$. Then as $\varphi$ was arbitrary $M,\bar{m}\equiv^{s}_{k}N,\bar{n}$.
Our second concept for generalisation is distance in a model and the associated local neighbourhoods. Classically this done via the Gaifman graph of a structure, the graph with vertices as elements of the model which are adjacent if they appear as part of a tuple in any modelled relation symbol. This lets use the standard distance metric on graphs to define a distance metric on any classical model in arbitrary signature. In our many-valued setting we can define a number of variants to the Gaifman graph metric; we start with a naïve version where we again link the the edges of the Gaifman graph to modelling. {defi} Let $A$ be a well-connected residuated lattice. We let $M$ be a $\mathcal{P}$ model and define the Gaifman graph $G(M)$ as the graph with vertex set $M$ and $mEn$ iff there is an $R\in\mathcal{P}$ and $\bar{s}\in M:m,n\in\bar{s}$ and $M\models R(\bar{s})$. We let $d(x,y)$ denote the standard distance metric on the graph $G(M)$, i.e. $d(m,n)$ is the length of the shortest path from $m$ to $n$ in $G(A,M)$.
For $\mathcal{P}$ relational without constant symbols, for $r\in\omega$ and $m\in M$ We define the $r$-sphere of $\bar{m}$ as:
$B(r,m):=\{n\in M:d(m,n)\leq r\},$
and use this interchangeably to denote the set of points and the induced strong substructure of $M$, i.e. where each $R\in\mathcal{P}$ is interpreted as the restriction of $R^{M}$ to $B(r,\bar{m})$. This naturally extends to use $B(r,\bar{m}),\bar{m}$ to denote the $\mathcal{P}\cup\{c_{1},...,c_{s}\}$-model where each $c_{i}$ is interpreted as $m_{i}$. For $\bar{m}\in M$ we define $B(r,\bar{m}):=\bigcup\limits_{m\in\bar{m}}B(r,m)$.
If $\mathcal{P}$ does contain constant symbols we require that all $r$-spheres for $m$ contain the ‘common core’ determined by the interpretation of the constant symbols.
$B(r,\varnothing)$ $:=\bigcup\limits_{c\in\mathcal{P}}\{n\in M:d(c^{M},n)\leq r\},$
$B(r,m)$ :=$:=\{n\in M:d(m,n)\leq r\}\cup B(r,\varnothing).$
Remark 12.
We note that for any $n,s\in B(r,m)$ $d(n,s)\leq 2r$.
One can directly define the $r$-sphere type of a tuple $\bar{m}$ in model $M$ to be the 00-isomorphism type of $B(r,\bar{m})$ that is $\varphi^{0}_{B(r,\bar{m})}(\bar{v})$. It then follows by Lemma 8 that for any other tuple $\bar{n}$ in a model $N$ $N^{\ast}\models\varphi^{0}_{B(r,\bar{m})}(B(r,\bar{n}))$ iff $B(r,\bar{m}),\bar{m}\cong B(r,\bar{n}),\bar{n})$ iff $M^{\ast}\models\varphi^{0}_{B(r,\bar{n})}(B(r,\bar{m}))$. Note however that strictly speaking the formula $\varphi^{0}_{B(r,\bar{m})}(\bar{v})$ depends on the exact listing of the elements of $B(r,\bar{m})$. This means we should really speak of an $r$-sphere type of $\bar{m}$. Given the primary purpose of discussing $r$-sphere types is to have a shorthand way of saying that local neighbourhoods are isomorphic, it proves more convenient to shift perspective and consider isomorphism types directly through isomorphism equivalence classes of finite models. {defi} Let $A$ be a well-connected residuated lattice. By an isomorphism type we mean an equivalence class for the isomorphism relation on $\mathcal{P}\cup\{c_{1},...,c_{s}\}$-models defined over $A$. We use the letters $\iota,\tau$ etc. to refer to isomorphism types.
If $\bar{m}\in M^{s}$ we say that $\bar{m}$ $r$-realises $\iota$, or equivalently that $\iota$ is the $r$-sphere type of $\bar{m}$ iff $B(r,\bar{m}),\bar{m}$ belongs to $\iota$.
Remark 13.
The direct via isomorphism definition aligns with syntactic definition of 00-isomorphism types in the sense that two tuples $\bar{m}$ and $\bar{n}$ have the same $r$-sphere type iff the $r$-sphere of $\bar{n}$ realises a formula $\varphi^{0}_{B(r,\bar{m})}(\bar{v})$ iff the $r$-sphere of $\bar{m}$ realises a formula $\varphi^{0}_{B(r,\bar{n})}(\bar{v}$.
The application of isomorphism types to spheres is enough to state Hanf locality. {thmC}[[Lib04]] Hanf Locality (Classical) Let $M,N$ be finite (classical) $\mathcal{P}$-models and $k\in\omega$. Letting $\langle r_{j}\rangle_{j\leq k}$ be recursively defined by $r_{0}=0$ $r_{j+1}=3r_{j}+1$. Let $e$ be the maximum size of the $r_{k}$-spheres in $M$ and $N$ and suppose that for each $r\leq r_{k}$ and sphere type $\iota$ either $i.$ or $ii.$ holds:
  1. i.
    $M$ and $N$ have the same number of elements that $r$-realise $\iota$.
  2. ii.
    Both $M$ and $N$ have more than $k\cdot e$ elements that $r$-realise $\iota$.
Then $M\equiv_{k}N$. This theorem fails in all relevant residuated lattices beyond the Boolean one. {exa} Let $\mathcal{P}$ be the language of weighted graphs, i.e. consist of a single binary relation symbol $E$. Let $A$ be any well-connected residuated lattice with at least $3$ elements. Let $a,b\in A\setminus\{1\}$ be distinct. We define the models $M$ and $N$ with the same two element domain $\{s,t\}$ with interpretation of $E$:
$E^{M}(s,s)=1$ $E^{M}(t,t)=1$ $E^{N}(s,s)=1$ $E^{N}(t,t)=1$
$E^{M}(s,t)=a$ $E^{M}(t,s)=a$ $E^{N}(s,t)=b$ $E^{N}(t,s)=a.$
Consider the sentence $\varphi:=\forall x\forall y(xEy\mathord{\reflectbox{/}}yEx)$. One can easily verify that $||\varphi||^{M}=1$ and $||\varphi||^{N}=(a\mathord{\reflectbox{/}}b)\land(b\mathord{\reflectbox{/}}a)<1$ (this is because $a\not=b$ and $x\mathord{\reflectbox{/}}y=1$ iff $x\leq y$ by the residuation property). Therefore $M\not\equiv_{2}N$.
By contrast $M$ and $N$ satisfy the assumption in the Hanf theorem. For $r\in\omega$ let $B_{M}(r,s)$ be the $r$-sphere of $s$ in $M$ and $B_{N}(r,s)$ be the $r$-sphere of $s$ in $N$, similarly for $t$. Now, $B_{M}(r,s)=\{s\}B_{N}(r,s)$ and $B_{M}(r,t)=\{t\}=B_{N}(r,t)$, and moreover the induced strong substructures are isomorphic given that $E^{M}$ and $E^{N}$ agree on $(s,s)$ and $(t,t)$. This means the $r$-sphere types of $s$ in $M$ and $N$ and $t$ in $M$ and $N$ are the same, and so for any sphere type $\iota$ $M$ and $N$ have the same number of elements that $r$-realise $\iota$.
There are natural alternatives to the modelling based Gaifman distance metric and this motivates the question of whether we might recover Hanf locality for some appropriate metric. {defi} Let $A$ be a well-connected residuated lattice and $a\in A$. For a $\mathcal{P}$-model we define its $a_{\geq}$-Gaifman graph as the graph with vertex set $M$ and $mEn$ iff there is an $R\in\mathcal{P}$ and $\bar{s}\in M:m,n\in\bar{s}$ and $||R(\bar{s})||^{M}\geq a$. We let $d_{\geq a}(x,y)$ denote the standard distance metric on the $a_{\geq}$-Gaifman graph of $M$. We define the corresponding $a$-sphere $B(r,\bar{m})$ and $a,r$-sphere type of $\bar{m}$ analogously.
We can repeat this again for the $a_{>}$ graph and distance metric defined $d_{>a}(x,y)$ with $>$ in place of $\geq$. We refer to these distance metrics as the (strict) threshold distance metrics.
[…]
Lemma 14.
Let $M,N$ be $\mathcal{P}$-models over a well-connected bounded residuated lattice $A$. Let $r\in\omega,\bar{m}\cup\{m\}\subseteq M$ and $\bar{n}\cup\{n\}\subseteq N$. Suppose that:
  1. i.
    $m\not\in B(2r+1,\bar{m})$,
  2. ii.
    $n\not\in B(2r+1,\bar{n})$,
  3. iii.
    $B(r,\bar{m})\cong B(r,\bar{n})$
  4. iv.
    $B(r,m)\cong B(r,n)$.
Then $B(r,\bar{m}m)\cong B(r,\bar{n}n)$.
Proof 3.4.
For the moment suppose that $\mathcal{P}$ contains no constant symbols. Letting $\pi$ and $\pi^{\prime}$ be the respective isomorphisms we consider the map $\pi^{+}:=\pi\cup\pi^{\prime}$. Conditions $i.$ and $ii.$ imply that $B(r,\bar{m})\cap B(r,m)=\varnothing=B(r,\bar{n})\cap B(r,n)$ so this is a well defined function and a bijection, it remains to check it is an isomorphism.
We first note that for any $s\in B(r,\bar{m})$ and $t\in B(r,m)$, if $d_{>\bot}(s,t)\leq 1$ then $d_{>\bot}(\bar{m},m)\leq 2r+1$ and $m\in B(2r+1,\bar{m})$ which is a contradiction. Therefore $d_{>\bot}(s,t)>1$ and in particular for any $s,t\in B(r,\bar{m}m)$ if $d_{>\bot}(s,t)\leq 1$ then either $s,t\in B(r,\bar{m})$ or $s,t\in B(r,m)$. Similarly, for any $s,t\in B(r,\bar{n}n)$ if $d_{>\bot}(s,t)\leq 1$ then either $s,t\in B(r,\bar{n})$ or $s,t\in B(r,n)$.
So now let $R\in\mathcal{P}$ and $\bar{s}\in B(r,\bar{m}m)$. We have three cases: $R^{M}(\bar{s})>\bot$, $R^{N}(\pi^{+}(\bar{s}))>\bot$ or both $R^{M}(\bar{s}=\bot=R^{N}(\pi^{+}(\bar{s}))$. Our check for the final case is immediate, so suppose $R^{M}(\bar{s})>\bot$. Then for all $i,j$ $d_{>\bot}(s_{i},s_{j})\leq 1$ and so either $\bar{s}\in B(r,\bar{m})$, and then
$R^{B(r,\bar{m}m)}(\bar{s})=R^{B(r,\bar{m})}(\bar{s})=R^{B(r,\bar{n})}(\pi(\bar{s}))=R^{B(r,\bar{n}n)}(\pi^{+}(\bar{s})),$
or $\bar{s}\in B(r,m)$ and similarly
$R^{B}(r,\bar{m}m)(\bar{s})=R^{B(r,m)}(\bar{s})=R^{B(r,n)}(\pi^{\prime}(\bar{s}))=R^{B(r,\bar{n}n)}(\pi^{+}(\bar{s}).$
If $R^{N}(\pi^{+}(\bar{s}))>0$ then for all $i,j$ $d_{>\bot}(\pi^{+}(s_{i}),\pi^{+}(s_{j}))\leq 1$ and either $\pi^{+}(\bar{s})\in B(r,\bar{n})$ or $\pi^{+}(\bar{s})\in B(r,n)$ and we can make the same identities.
Now we consider the case where we allow constant symbols. In this case, $B(r,\bar{m})\cap B(r,m)=B(r,\varnothing)$. The potential problem with combining our isomorphisms is they may conflict on the common core. Notably however, the assumption that $m\not\in B(2r+1,\bar{m})$ further implies that $\{t\in M:d_{>\bot}(t,m)\leq r\}\cap B(r,\varnothing)=\varnothing$ and thus we may ’break’ in favour of the isomorphism for $B(r,\bar{m})$. More precisely, we let $\pi$ be the isomorphism between $B(r,\bar{m})$ and $B(r,\bar{n})$, $\pi^{\prime}$ be the restriction of the isomorphism between $B(r,m)$ and $B(r,n)$ onto the sets $\{t\in M:d_{>\bot}(t,m)\leq r\}$ and $\{t\in N:d_{>\bot}(t,n)\leq r\}$ and consider the map $\pi^{+}=\pi\cup\pi^{\prime}$. This is a well defined bijection due to acting on disjoint sets and we again must check it is an isomorphism. For constant symbols this is immediate as $\pi^{+}(c^{B(r,\bar{m}m})=\pi(c^{B(r,\bar{m})})=c^{B(r,\bar{n})}=c^{B(r,\bar{n}n)}$, and for relation symbols this is just as before except we work with the pairs of disjoint sets $B(r,\bar{m})$, $\{t\in M:d_{>\bot}(t,m)\leq r\}$ and $B(r,\bar{n})$, $\{t:\in N:d_{>\bot}(t,n)\leq r\}$.
Theorem 15 (Hanf’s Theorem for Residuated Lattices).
Let $M,N$ be finite $\mathcal{P}$-models over a well-connected residuated bounded lattice $A$, $d$ be the strict $\bot$-threshold distance and $k\in\omega$. We define the sequence $\langle r_{j}\rangle_{j\leq k}$ recursively by $r_{0}=0$ and $r_{j+1}=3r_{j}+1$. Let $e$ be the maximum size of the $r_{k}$-sphere in $M$ and $N$ and suppose that for each $r\leq r_{k}$ and sphere type $\iota$ either $i.$ or $ii.$ holds:
  1. i.
    $M$ and $N$ have the same number of elements that $r$-realise $\iota$.
  2. ii.
    Both $M$ and $N$ have more than $k\cdot e$ elements that $r$-realise $\iota$.
Then $M\equiv^{s}_{k}N$.
Proof 3.5.
We are able to essentially follow the textbook proof for the classical case (e.g. Ebbinghaus and Flum [EF95, Theorem 1.4.1] or [Lib04, Theorem 4.24]. We show that $\langle I_{j}\rangle_{j\leq k}$ is a b&f-system from $M$ and $N$ where we define:
$I_{j}:=\{\bar{m}\mapsto\bar{n}\text{ p.iso for }M,N:(B(r_{j},\bar{m}),\bar{m})\cong(B(r_{j},\bar{n}),\bar{n})\text{ and }|\bar{m}|\leq k-j\}.$
For $j=k$ we have $\bar{m}=\varnothing$ and take $I_{k}=\{\varnothing\mapsto\varnothing\}$ as a trivial partial isomorphism. This still requires checking that $B(r_{k},\varnothing)$ as defined in $M$ is isomorphic to $B(r_{k},\varnothing)$ as defined in $N$. This follows easily by our assumption, if we let any $m\in M$ and $\iota$ be its $r_{k}$-sphere type, $\exists n\in N$ with $r_{k}$-sphere type $\iota$ and so $B(r_{k},m)\cong B(r_{k},n)$. These spheres contain the local neighbourhoods of the interpreted constant symbols in their respective models, so we can restrict our isomorphism into those neighbourhoods.
Now we turn to the back and forth conditions. By symmetry it is enough to prove the forth condition. Accordingly, let $0\leq j<k,m\in M$ and $\bar{m}\mapsto\bar{n}\in I_{j+1}$ witnessed by say $\pi$.
Case $1$: $m\in B(2r_{j}+1,\bar{m})$. In this case $B(r_{j},\bar{m}m)\subseteq B(r_{j+1},\bar{m})$, as for any $t\in B(r_{j},\bar{m}m)$ either $d(t,m_{i})\leq r_{j}\leq r_{j+1}$ or $d(t,m)\leq r_{j}$ and then there is some $i:d(t,m_{i})\leq d(t,m)+d(m,m_{i})\leq r_{j}+2r_{j}+1=3r_{j}+1=r_{j+1}$). Therefore we can consider $n=\pi(m)$ and $\pi$ restricts to an isomorphism on $B(r_{j},\bar{m}m)\cong B(r_{j},\bar{n}n)$. Then $\bar{m}m\mapsto\bar{n}n\in I_{j}$ as required.
Case $2$: $m\not\in B(2r_{j}+1,\bar{m})$. We let $\iota$ be the $r_{j}$-sphere type of $m$. By the existence of $\pi$ we claim that $B(2r_{j}+1,\bar{m})$ and $B(2r_{j}+1,\bar{n})$ have the same number of elements of $r_{j}$-sphere type $\iota$. For any $t\in B(2r_{j}+1,\bar{m})$ $B(r_{j},t)\subseteq B(2r_{j}+1,\bar{m})\subseteq B(r_{j+1},\bar{m})$ and therefore $(B(r_{j},s),s)\cong\pi[B(r_{j},s),s]=B(r_{j},\pi(s),\pi(s))$. Thus $s$ and $\pi(s)$ have the same $r_{j}$-sphere type and the injectivity of $\pi$ implies the number of elements of $r_{j}$-sphere type $\iota$ in $B(2r_{j}+1,\bar{m})$ is less than or equal to that number in $B(2r_{j},\bar{n})$. By symmetry with $\pi^{-1}$ we have equality. Now, if $i$. holds for $\iota$ as $m\in M\setminus B(2r_{j}+1,\bar{m})$ and is of $r_{j}$-sphere type $\iota$ there must exists an $n\in N\setminus B(2r_{j}+1,\bar{n})$ of $r_{j}$-sphere type $\iota$ otherwise.
In the case that $ii.$ holds for $\iota$, we note that the size of $B(2r_{j}+1,\bar{m})$ is $\leq r_{k}\leq e$ by assumption and so $|B(2r_{j},m)|\leq$ length$(\bar{a})\cdot e\leq k\cdot e$. In particular, the number of elements of $r_{j}$-sphere type $\iota$ in $B(2r_{j},\bar{m})$ and $B(2r_{j}+1,\bar{n})$ is at most $k\cdot e$ and therefore $\exists n\in N\setminus B(2r_{j}+1,\bar{n})$ of $r_{j}$-sphere type $\iota$.
In either case we have found $n\not\in B(2r_{j}+1,\bar{n})$ with $B(r_{j},m)\cong B(r_{j},n)$ and we may therefore apply our previous lemma and conclude that $B(r_{j},\bar{m}m),\bar{m}m\cong B(r_{j},\bar{n}n),\bar{n}n$ witnessing that $\bar{m}m\mapsto\bar{n}n\in I_{j}$.
The fact we can follow the classical proof of Hanf’s theorem essentially unmodified emphasises that the real interest is in the condition within the theorem, a thought we return to in Section 6. As an immediate first step we note that it implies another form of locality, due to Fagin, Stockmeyer and Vardi [FSV94]. {defi} Let $A$ be a well-connected residuated lattice. Let $M,N$ be finite $\mathcal{P}$-models over $A$. We write $M\leftrightarrows_{r}N$ iff for each isomorphism type $\tau$ of $\mathcal{P}\cup\{c\}$-structures the number of elements of $M$ and $N$ with $r$-sphere type $\tau$ is the same.
Remark 16.
By remark 13 this condition is equivalent to there being a bijection $f\colon M\rightarrow N$ such that for all $m\in M$ $(B(r,\bar{m}),\bar{m})\cong B(r,f(\bar{m})),f(\bar{m})$.
Corollary 17.
Let $A$ be a well-connected residuated lattice. Let $M,N$ be finite $\mathcal{P}$-models over $A$ and $k\in\omega$. Suppose $M,\bar{m}\leftrightarrows_{3^{k}}N,\bar{n}$. Then $M,\bar{m}\equiv^{s}_{k}N,\bar{n}$.
Proof 3.6.
We aim to show $M,\bar{m}$ and $N,\bar{n}$ satisfy the Hanf theorem criterion. We already have the condition for $3^{k}$ so we just need to check that if $M,\bar{m}$ and $N,\bar{n}$ have the same number of elements of $r$-sphere type $\tau$ then they also have the same number of elements of $l$-sphere type $\iota$ for all $l\leq r$. But this is clear; two elements have the same $r$-sphere type iff $B(r,x)\cong B(r,y)$ and have the same $l$-sphere type iff $B(l,x)\cong B(l,y)$ and two spheres being isomorphic certainly implies smaller spheres are.
4. Gaifman Locality
We now want to follow the pattern of Hanf locality with Gaifman locality; set up the many-valued counterparts to the key concepts the proof and demonstrate the resulting theorem goes through as it does classically. There are however some important differences. Firstly, we will need to recover some additional syntactic machinery which puts us into a more restricted setting than our quite general one for Hanf. We will introduce these restrictions alongside the machinery they are required for. Secondly, we do not recover the usual classical Gaifman theorem. In the classical setting, a common proof of Gaifman’s theorem (for both finite and infinite models) proceeds via a main lemma to which the theorem comes as a corollary [EF95, Theorem 1.5.1]. This implication relies on compactness. {propC}[[EF95]] Let $\Phi$ be a set of first-order classical sentences. Assume that any two structures that satisfy the same sentences of $\Phi$ are elementary equivalent. Then any first-order sentence is equivalent to a boolean combination of sentences of $\Phi$.
Proof 4.1 (Proof sketch).
Take an arbitrary (classical) structure $A$ and consider
$\Phi(A):=\{\psi\in\Phi:A\models\psi\}\cup\{\neg\psi:\psi\in\Phi,A\models\neg\psi\}.$
An easy application of compactness tells us that if $\varphi\Relbar\joinrel\mathrel{|}\models\bigvee\limits_{A\models\varphi}\bigwedge\Phi(A)$ then there is a finite collection of structures $A_{i}$ and finite $\Phi^{i}\subseteq\Phi(A_{i})$ $\varphi$ is equivalent to the sentence:
$\psi:=\bigvee\limits_{A_{i}}\bigwedge\Phi^{i}.$
The status of compactness for non-classical models is a significant area of research in its own right. As a brief mention of some positive results, compactness has been recovered for models defined over a fixed finite MTL-algebra [Del13, Theorem 4.4] and models defined over the standard $[0,1]$ based Łukasiewicz and Gödel algebras, but fails for the $[0,1]$ product algebra [H0́2]. Here our restriction to finite models proves significant. Just as compactness fails for finite classical models [EF95, Chapter 0] the same formulas witness its failure for finite models defined over non-trivial, well-connected residuated lattices. It is a proof of the hypothesis in this corollary that we generalise into the abstract setting, and which we refer to as Gaifman’s Lemma. This is still of significant interest, as we will see later it is enough to recover the application to definable queries.
Theorem 18 (Gaifman’s Lemma (Classical)).
Suppose $A$ and $B$ satisfy the same basic local sentences. Then $A\equiv B$.
Basic local sentences are a particular syntactic form based on local formula and it is this we need to form an abstracted version of. This is enabled by another syntactic encoding, this time of the distance metric. We define the formula $\theta_{r}(x,y)$ inductively:
$\theta_{0}(x,y)$ $:=x=y$
$\theta_{1}(x,y)$ $:=\bigvee\limits_{R\in\mathcal{P}}\exists u_{1}...\exists u_{a(R)}[R(u_{1},...,u_{a(R)})\land\bigvee\limits_{1\leq i,j\leq a(R)}(u_{i}=x\land u_{j}=y)]$
$\theta_{r+1}(x,y)$ $:=\exists z(d_{k}(x,z)\land d_{1}(z,y)).$
These same formulas applied to finite models over a well-connected residuated lattice $A$ capture a refined measure of distance. For any $\mathcal{P}$-model over $A$ we define the Gaifman matrix $G_{v}(M)$ as the $|M|\times|M|$ matrix with components:
$v_{m,n}:=sup\{R^{M}(\bar{c})\in A:R\in\mathcal{P}m,n\in\bar{c}\},$
that is the supremum of values taken by atomic formula in which $m$ and $n$ occur.
A path through $G_{v}(A,M)$ is a sequence of elements $M$ and for a path of length $r-1$, its label is the $r$-sequence $\langle v_{m_{i},n_{i}}\rangle_{i\leq r}$ and its weight is $inf\{v_{m_{i},n_{i}}\in A:1\leq i\leq r\}$. It is straightforward to verify that $||\theta_{r}(m,n)||^{M}$ is the supremum of the weight of paths through $G_{v}(M)$ from $m$ to $n$ of length at most $r$ and as corollary to this that it encodes the distance in the naïve Gaifman graph.
Corollary 19.
Let $A$ be a well-connected residuated lattice, $M$ a finite $\mathcal{P}$-model over $A$, $m,n\in M$ and $r\in\omega$.
$M\models\theta_{r}(m,n)\text{ iff }||\theta_{r}(m,n)||^{M}\geq 1\text{ iff }d(m,n)\leq r\text{ iff }n\in B_{a}(r,m).$
Remark 20.
Note that we only ask for our distance formulas to encode distance at the level of the modelling relation. [^6] This contrasts with the approach of Bizière, Grädel and Naaf who require an encoding of distance that allows for quantification within distance to take the actual value. This difference is driven by the needs of the respective investigations which we will return to in Section 5.
This corollary relies on our limited form of witnessing applying to both the quantifiers $\exists$/$\forall$ / ∀ and the lattice connectives $\lor$/$\land$ / ∧. Given any $a\in A$, if this limited form of witnessing holds at $a$ then we can encode the threshold distance metrics $d_{a}(x,y)$ for the $a$-Gaifman graph using the standard expansion of a model. In particular, when $M$ is finite and $A$ is a chain this applies to all $a\in A$.
Corollary 21.
Let $A$ be a residuated chain, $a\in A$, $M$ a finite $\mathcal{P}$-model over $A$, $m,n\in M$ and $r\in\omega$.
$M^{\ast}\models a\mathord{\reflectbox{/}}\theta_{r}(m,n)\text{ iff }||\theta_{r}(m,n)||^{M}\geq a\text{ iff }d_{a}(m,n)\leq r\text{ iff }n\in B_{a}(r,m).$
The classical distance formulas naturally come paired with their negations, with the property that $M\models\neg\theta_{r}(m,n)$ iff $d(m,n)>M$, and these are essential to the definition of local formula/basic local sentences. In general residuated lattices do not have such a classical-like negation term, but we can consider a still interesting smaller class which do and consequently for which a ‘far’ distance formulas can be defined, namely those $A$ with a co-atom. [^7] The name is taken from the integral case where $1=\top$ and c is a co-atom in the familiar sense. An element $c\in A$ is a co-atom iff $\forall b\in A$ $1\not\leq b$ iff $b\leq c$. The co-atom lets us define a classical like negation term $t(x):=x\mathord{\reflectbox{/}}c$ in the language $P^{A}$ where $M^{\ast}\models\varphi$ iff $M^{\ast}\not\models t(\varphi)$. This lets us define the distance formulas as we did classically namely:
$\rho_{r}(x,y)=t(\delta_{r}(x,y)).$
In the case of the strict threshold distance metrics the situation is reversed, the standard expansion of a model can directly express the $\rho$ distance formula and we use the negation like term to express the $\delta$ formula.
[…]
It is worth noting that for threshold/strict threshold distance we could equally use an element that acts like a co-atom/atom ($\forall b\in A$ $a\not<b$ iff $c\leq b$) for $a$ itself, directly defining $\rho_{r}(x,y):=\theta_{r}(m,n)\mathord{\reflectbox{/}}c$ or $\delta_{r}(x,y):=c\mathord{\reflectbox{/}}\theta_{r}(m,N)$ respectively. We will default to the assumption of the $1$ co-atom, but the work could be equally completed using the alternative definitions of $\delta_{r}$ and $\rho_{r}$. To that end, from now on we proceed for an arbitrary (strict) threshold distance metric and use $\delta_{r}(x,y)$ and $\rho_{r}(x,y)$ as shorthand for any appropriate encoding of distance, i.e. formulas such that
$M^{\ast}\models\delta_{r}(m,n)\text{ iff }d(m,n)\leq r,\,M^{\ast}\models\rho_{r}(m,n)\text{ iff }d(m,n)>r.$
Classically, the encoding of distance allows one to associate to any formula $\varphi(\bar{x},\bar{y})$ a formula $\varphi^{r,\bar{x}}(\bar{x},\bar{y})$ such that for any classical model $M$, $\bar{m}\in M$ and $\bar{n}\in B(k,\bar{m})$:
$M\models\varphi^{r,\bar{x}}(\bar{m},\bar{n})\text{ iff }B(r,\bar{m})\models\varphi(\bar{m},\bar{n}).$
[…]
Proposition 22.
Let $\mathcal{L}$ be an algebraic signature and $\mathcal{P}$ a predicate language. Let $A$ be a $\mathcal{L}$-algebra and $M$ a $\mathcal{P}$-model over $A$. Let $\varphi,\lambda_{1},...,\lambda_{n}$ be $\mathcal{P}$-formulas where $x$ is not free in any $\lambda_{i}$ and $\circ\in\mathcal{L}$ be any $n+1$-ary connective.
If $M$ is a witnessed $\mathcal{P}$-model then:
  1. a)
    $||\forall x(\circ(\varphi,\lambda_{1},...,\lambda_{n})||^{M}\leq||\circ(\exists x\varphi,\lambda_{1},...,\lambda_{n})||^{M}\leq||\exists x(\circ(\varphi,\lambda_{1},...,\lambda_{n}))||^{M}$,
  2. b)
    $||\forall x(\circ(\varphi\lambda_{1},...,\lambda_{n}))||^{M}\leq||\circ(\forall x\varphi,\lambda_{1},...,\lambda_{n})||^{M}\leq||\exists x(\circ(\varphi,\lambda_{1},...,\lambda_{n})||^{M}$,
If $\circ$ is monotone for the position of $\varphi$ we have:
  1. c)
    $||\exists x(\circ(\varphi,\lambda_{1},...,\lambda_{n}))||^{M}\leq||\circ(\exists x\varphi,\lambda_{1},...,\lambda_{n})||^{M}$,
  2. d)
    $||\circ(\forall x\varphi,\lambda_{1},...,\lambda_{n})||^{M}\leq||\forall x(\circ(\varphi,\lambda_{1},...,\lambda_{n}))||^{M}$,
When $\circ$ is antitone for the position of $\varphi$ we have:
  1. e)
    $||\exists x(\circ(\varphi,\lambda_{1},...,\lambda_{n}))||^{M}\leq||\circ(\forall x\varphi,\lambda_{1},...,\lambda_{n})||^{M}$,
  2. f)
    $||\circ(\exists x\varphi,\lambda_{1},...,\lambda_{n})||^{M}\leq||\forall x(\circ(\varphi,\lambda_{1},...,\lambda_{n})||^{M}$.
Consequently, if every position of each connective $\circ\in\mathcal{L}$ is monotone or antitone then for every formula $\varphi$ there is a formula $\psi$ in prenex normal form, i.e. a string of quantifiers applied to a quantifier-free formula that is strongly equivalent to $\varphi$, i.e. such that for any finite $\mathcal{P}$ model $M$ and $\bar{m}\in\mathcal{P}$
$||\varphi(\bar{m})||^{M}=||\psi(\bar{m})||^{M}$
Proof 4.2.
a) By $\exists$-witnessing
$||\circ(\exists x\varphi,\lambda_{1},...,\lambda_{n})||=\circ^{A}(||\exists x\varphi||,||\lambda_{1}||,...,\lambda_{n}||)=\circ^{A}(||\varphi(m)||,||\lambda_{1}||,...,\lambda_{n}||)$
for some $m\in M$.
$\therefore||\forall x(\circ(\varphi,\lambda_{1},...,\lambda_{n})||^{M}=inf_{m}\{\circ^{A}(||\varphi(m)||,||\lambda_{1}||,...,||\lambda_{n}||)\}\\ \leq||\circ(\exists x\varphi,\lambda_{1},...,\lambda_{n})||\leq sup_{c}\{\circ^{A}(||\varphi(m),\lambda_{1},...,\lambda_{n})||)\}=||\exists x(\circ(\varphi,\lambda_{1},...,\lambda_{n}))||$
The proof for b) is symmetric.
c) Letting $m\in M$, $\circ$ monotone for $\varphi$’s position implies:
$||\circ(\varphi(m),\lambda_{1},...,\lambda_{n})||=\circ^{A}(||\varphi(m)||,||\lambda_{1}||,...,||\lambda_{n}||)\\ \leq\circ^{A}(||\exists x\varphi||,||\lambda_{1}||,...,||\lambda_{n}||)=||\circ(\exists x\varphi,\lambda_{1},..,\lambda_{n})||.$
Again d) is symmetric.
e) Let $m\in M$. Then as $\circ$ is antitone for $\varphi$’s position,
$||\circ^{A}(||\varphi(m)||,||\lambda_{1}||,...,||\lambda_{n}||)\leq||\circ^{A}(||\forall x\varphi||,||\lambda_{1}||,...,||\lambda_{n}||)=||\circ(\forall x\varphi,\lambda_{1},...,\lambda_{n})||.$
As $m\in M$ was arbitrary
$||\exists x(\circ(\varphi,\lambda_{1},...,\lambda_{n}))||\leq||\circ(\forall x\varphi,\lambda_{1},...,\lambda_{n}))||.$
Again f) is symmetric.
The conclusion for prenex normal form is then a straightforward induction utilising each of the quantifier shifts and renaming of variables as appropriate.
We define our relative quantified formulas for formulas in prenex normal form, and factor through these for formulas not in prenex normal form. {defi} Let $\varphi(\bar{x},\bar{y})$ be a formula in prenex normal form. We define its relativisation to $B(r,\bar{m})$ using the distance formulas.
$[\exists z\psi(\bar{x},\bar{y},z)]^{r,\bar{x}}$ $:=\exists z(\delta_{r}(\bar{x},z)\land\psi^{r,\bar{x}}(\bar{x},\bar{y},z)).$
$[\forall z\psi(\bar{x},\bar{y},z)]^{r,\bar{x}}$ $:=\forall z(\rho_{r}(\bar{x},z)\lor\psi^{r,\bar{x}}(\bar{x},\bar{y},z)).$
Letting $\varphi(\bar{x},\bar{y})$ be any $\mathcal{P}$-formula we define its relativisation through a strongly equivalent formula in prenex normal form.
$\varphi^{r,\bar{x}}(\bar{x},\bar{y})=[Q_{1}z_{1}...Q_{n}z_{n}\psi(\bar{x},\bar{y},\bar{z})]^{r,\bar{x}}$
Lemma 23.
Let $A$ be a residuated chain with co-atom and consider an encodable distance metric for finite $\mathcal{P}$-models over $A$. Letting $M$ be a finite $\mathcal{P}$-model over $A$, for any $\mathcal{P}^{\ast}$-formula $\varphi(\bar{x},\bar{y})$ any $\bar{m}\in M$, $r\in\omega$, $\bar{n}\in B(r,\bar{m})$ we have:
$M^{\ast}\models\varphi^{(r,\bar{m})}(\bar{m},\bar{n})\text{ iff }B(r,\bar{m})^{\ast}\models\varphi(\bar{m},\bar{n}).$
Proof 4.3.
It is enough to prove the lemma for $\varphi$ in prenex normal form; we proceed by induction on the number of quantifiers in $\varphi$. When $\varphi$ is quantifier free $\varphi^{k,\bar{x}}(\bar{x},\bar{y})=\varphi(\bar{x},\bar{y})$ and so for $\bar{n}\in B(k,\bar{m})$
$||\varphi^{r,\bar{x}}(\bar{m},\bar{n})||^{M}=||\varphi(\bar{m},\bar{n})||^{M}=||\varphi(\bar{m},\bar{n})||^{B(k,\bar{m})}.$
So suppose for $l\in\omega$ that for all formulas $\psi$ in prenex normal form with at most $l$ quantifiers that
$M^{\ast}\models[\varphi(\bar{m},\bar{n})]^{k,\bar{x}}\text{ iff }B(k,\bar{m})^{\ast}\models\varphi(\bar{m},\bar{n}).$
We consider $QzQ_{1}z_{1}...Q_{l}z_{l}(\varphi(\bar{x},\bar{z},\bar{y}))$. Either $Q=\exists$ or $\forall$. The two cases are symmetric, we detail the $Q=\exists$ case.
$[\exists z...Q_{l}z_{l}(\varphi(\bar{x},\bar{z},\bar{y}))]^{k,\bar{x}}=\exists z(Q_{1}z_{1}...Q_{l}z_{l}\varphi^{r,\bar{x}}(\bar{x},\bar{z},\bar{y},z)\land\delta_{k}(\bar{x},z)).$
Suppose $M^{\ast}$ models this formula. Then,
$1\leq||\exists z(Q_{1}z_{1}...Q_{l}z_{l}\varphi^{r,\bar{x}}(\bar{m},\bar{z},\bar{n},z)\land\delta_{k}(\bar{x},z))||^{M}=||Q_{1}z_{1}...Q_{l}z_{l}\varphi^{r,\bar{x}}(\bar{m},\bar{z},\bar{n},n)\land\delta_{k}(\bar{m},z)||^{M}$
for some $n\in M$ by witnessing. This implies $||Q_{1}z_{1}...Q_{l}z_{l}\varphi^{r,\bar{x}}(\bar{m},\bar{z},\bar{n},n)||^{M}$ and $||\delta_{k}(\bar{m},n)||^{M}$ are both $\geq 1$, in particular $n\in B(k,\bar{m})$. By the inductive hypothesis:
$||Q_{1}z_{1}...Q_{l}z_{l}\varphi(\bar{m},\bar{z},\bar{n},n))||^{B(k,\bar{m})}\geq 1\text{ and }n\in B(k,\bar{m})$
[…]
Conversely, if $||\exists z(Q_{1}z_{1}...Q_{l}z_{l}\varphi(\bar{m},\bar{z},\bar{n},z))||^{B(k,\bar{m})}\geq 1$, then by witnessing we have $n\in B(k,\bar{m})$, i.e. $n\in M:||\delta_{k}(\bar{m},n)||\geq 1$ where
$||Q_{1}z_{1}...Q_{l}z_{l}\varphi(\bar{z},\bar{n},n))||^{B(k,\bar{m})}\geq 1.$
By induction hypothesis $||Q_{1}z_{1}...Q_{l}z_{l}\varphi^{r,\bar{x}}(\bar{m},\bar{z},\bar{n},z))||^{M}\geq 1$ which implies,
$||\exists zQ_{1}z_{1}...Q_{l}z_{l}\varphi^{k,\bar{m}}(\bar{m},\bar{z},\bar{n},z))\land\delta_{k}(\bar{m},n)||^{M}=||[\exists z(Q_{1}z_{1}...Q_{l}z_{l}(\bar{m},\bar{z},\bar{n},z))]^{k,\bar{x}}||^{M}\geq 1.$
In order to guarantee that all models considered are witnessed we therefore restrict $A$ to be linear i.e. a residuated chain. We can now define our local formula and basic local sentences. {defi} Let $A$ be a residuated chain and consider the distance formulas $\delta_{r}(x,y)$, $\rho_{r}(x,y)$ for any of the considered distance metrics. We define two special classes of formulas.
The local formulas are exactly those of the form $\varphi^{r,\bar{x}}(\bar{x})$ with no free variables beyond those used to define the neighbourhood quantifiers. We call a sentence basic local iff it has the form
$\exists x_{1}...\exists x_{n}\bigwedge\limits_{1\leq i,j\leq n}(\rho_{2r}(x_{i},x_{j})\land\bigwedge\limits_{1\leq i\leq n}\psi^{r,x_{i}}(x_{i}).$
We are now able to state and prove our version of Gaifman’s Lemma.
Theorem 24 (Gaifman Lemma for Residuated Chains).
Let $\mathcal{P}$ be a finite relational language and let $M,N$ be finite $\mathcal{P}$-models over a residuated chain $A$ with co-atom. Consider any encodable distance metric for $\mathcal{P}$-models with distance formulas $\delta_{r}(x,y)$ and $\rho_{r}(x,y)$.
Let $k\in\omega$, $\bar{m}\in M$ and $\bar{n}\in N$ and suppose that:
  1. i.
    For all local $\mathcal{P}$-formulas $\varphi^{7^{k},\bar{x}}(\bar{x})$ $M^{\ast}\models\varphi^{7^{k},\bar{x}}(\bar{m})$ iff $N^{\ast}\models\varphi^{7^{k},\bar{x}}(\bar{n})$.
  2. ii.
    $M^{\ast}$ and $N^{\ast}$ satisfy the same basic local sentences.
Then $M,\bar{m}\equiv^{s}_{k}N,\bar{n}$.
Proof 4.4.
Once again we are able to mirror the textbook classical proof [EF95, Theorem 1.5.1]. The addition of a leading sequences requiring some minor initial adjustments. We show that $\langle I_{j}\rangle_{j\leq k}$ is a b&f-system for $M^{\ast}$ and $N^{\ast}$ where we define:
$I_{j}:=\{\bar{m}\bar{u}\mapsto\bar{n}\bar{v}\text{ p.iso}:|\bar{u}|\leq k-j\text{ and }(B(7^{j},\bar{m}\bar{u}),\bar{m}\bar{u})^{\ast}\equiv^{s}_{g(j)}(B(7^{j},\bar{n}\bar{v}),\bar{n}\bar{v})^{\ast}\}.$
where $g(j)$ is a sufficiently increasing function defined by constraints established in the course of the proof. The conclusion then follows by corollary 11. To ease notation we will simply write $\bar{u}\mapsto\bar{v}$ for a p.iso and assume the leading $\bar{m}\mapsto\bar{n}$ prefix.
We first claim that $\bar{m}\mapsto\bar{n}\in I_{m}$. Letting $\varphi(\bar{x})$ be any $\mathcal{P}^{\ast}\cup\{\bar{x}\}$-sentence, we observe:
$B(7^{j},\bar{m})^{\ast},\bar{m}\models\varphi(\bar{m})\text{ iff }M^{\ast}\models\varphi^{7^{j},\bar{x}}(\bar{m})\text{ iff }N^{\ast}\models\varphi^{7^{j},\bar{x}}(\bar{n})\text{ iff }B(7^{j},\bar{n})^{\ast},\bar{n}\models\varphi(\bar{n}),$
where the first and third bi-conditional holds by Lemma 23 and the middle bi-conditional from assumption $i.$.
By symmetry we just need to show the forth property holds for the system, that is for $0\leq j\leq k$, $u\in M$ and $\bar{u}\mapsto\bar{v}\in I_{j+1}$ we want to find $v\in N$ such that $\bar{u}u\mapsto\bar{v}v$ is a p.iso and $(B(7^{j},\bar{u}u),\bar{u}u)^{\ast}\equiv^{s}_{g(j)}(B(7^{j},\bar{v}v),\bar{v}v)^{\ast}$. We introduce a useful abbreviations for $\bar{d}\in D$:
$\psi^{j}_{\bar{d}}(\bar{x}):=[\varphi^{g(j)}_{B(7^{j},\bar{d}),\bar{d}}(\bar{x})]^{7^{j},\bar{x}}.$
Combing lemmas 23 and corollary 11 yields the important behaviour of the formula $\psi^{j}_{\bar{d}}(\bar{x})$ for any model $S$ and $\bar{s}\in S$:
$S^{\ast}\models\psi^{j}_{\bar{d}}(\bar{s})\text{ iff }B(7^{j},\bar{s})^{\ast}\models\varphi^{g(j)}_{B(7^{j},\bar{d}),\bar{d})}(\bar{s})\text{ iff }B(7^{j},\bar{s})^{\ast},\bar{s}\equiv^{s}_{g(j)}B(7^{j},\bar{d})^{\ast},\bar{d}.$
Case $1$: $u\in B(2\cdot 7^{j},\bar{u})$. Then $u\in B(7^{j+1},\bar{u})$ and $B(7^{j+1},\bar{u})^{\ast}\models\exists z(\delta_{a}(\bar{u},z)\leq 2\cdot 7^{j}\land\psi^{j}_{\bar{u}u}(\bar{u}z))$. Forcing $g(j+1)$ large enough such that the quantifier rank of this formula is within $g(j+1)$ we get by the inductive assumption:
$B(7^{j+1},\bar{v})^{\ast},\bar{v}\models\exists z(\delta_{2\cdot 7^{j}}(\bar{v},z)\land\psi^{j}_{\bar{u}u}(\bar{v}z)),$
and by witnessing we get $c\in B(2\cdot 7^{j}),\bar{v})$ such that $B(7^{j+1},\bar{v})$ such that
$B(7^{j+1},\bar{v})^{\ast},\bar{v}\models\psi^{j}_{\bar{u}u}(\bar{v}v)\text{ i.e. }B(7^{j},\bar{u}u)^{\ast},\bar{u}u\equiv^{s}_{g(j)}B(7^{j},\bar{v}v)^{\ast},\bar{v}v,$
therefore $\bar{u}u\mapsto\bar{v}v\in I_{j}$.
Case $2$: $u\not\in B(2\cdot 7^{j},\bar{u})$, that is $B(7^{j},\bar{u})\cap B(7^{j},u)=\varnothing$. For $s\geq 1$ we define formula $\gamma_{s}(\bar{x})$ which expresses that $\{\bar{x}\}$ is a $4\cdot 7^{j}$-scattered subset where every members’ $7^{j}$ $a$-sphere has the same $g(j)$ isomorphism type as element of a model $u$.
$\gamma_{s}:=\bigwedge_{1\leq r,t\leq s}\rho_{4\cdot 7^{j}}(x_{r},x_{t})\land\bigwedge\limits_{1\leq r\leq s}\psi^{j}_{y}(x_{r}).$
We note that for $\bar{d}\in D$ $D^{\ast}\models\gamma_{s}(\bar{d})$ iff for all $1\leq r,t\leq s$ $D^{\ast}\models\rho_{4\cdot 7^{j}}(d_{r},d_{t})$ and $D^{\ast}\models\psi^{j}_{a}(d_{r})$ iff $\{\bar{d}\}$ forms a $4\cdot 7^{j}$-scattered subset of $D$ and for each $d\in\bar{d}$ $B(7^{j},u)^{\ast},u\equiv^{s}_{g(j)}B(7^{j},d)^{\ast},d$ iff $\{\bar{d}\}$ form a $4\cdot 7^{j}$-scattered subset and for each $d\in\bar{d}$ the $7^{j}$ $a$-sphere of $d$ has the same $g(j)$ isomorphism type as $u$.
We compare the relative cardinalities, defined as $e$ and $i$ of, maximal subsets satisfying this in $M^{\ast}$ and $B(2\cdot 7^{j},\bar{u})^{\ast}$ respectively. That is, we define $e$ and $i$ such that the following formulas hold:
$a.\,B(7^{j+1},\bar{u})^{\ast}\models\exists x_{1}...\exists x_{e}(\bigwedge\limits_{1\leq r\leq e}\delta_{2\cdot 4^{j}}(\bar{u},x_{r})\land\gamma_{e})$
$b.\,B(7^{j+1},\bar{u})^{\ast}\not\models\exists x_{1}...\exists x_{e+1}(\bigwedge\limits_{1\leq r\leq e+1}\delta_{2\cdot 4^{j}}(\bar{u},x_{r})\land\gamma_{e+1})$
$c.\,M^{\ast}\models\exists x_{1}...\exists x_{i}\gamma_{i},\qquad d.\,M^{\ast}\not\models\exists x_{1}...\exists x_{i+1}\gamma_{i+1}$
An easy consequence of witnessing establishes these represent the appropriate cardinalities. If no $i$ exists we set $i=\omega$; by contrast $e$ is bounded by the length of $\bar{u}$ (i.e. $k+|\bar{m}|$) since any two elements in the same sphere of radius $2\cdot 7^{j}$ have a distance at most $4\cdot 7^{j}$. Additionally, $e\leq i$ and claim that $e$ and $i$ as determined in $N^{\ast}$ and $B(7^{j+1},\bar{v})$ are the same. In the case of $N^{\ast}$ this is because the sentences $\exists x_{1}...\exists x_{i}\gamma_{i}$ are basic $a$-local and so $M$ and $N$ are in mutual agreement to their validity. For $B(7^{j+1},\bar{v})^{\ast}$ we recall from the forth hypothesis that $(B(7^{j+1},\bar{u})^{\ast},\bar{u})\equiv^{s}_{g(j+1)}(B(7^{j+1},\bar{v})^{\ast},\bar{v})$ so provided $g(j+1)$ is greater than the quantifier rank of other above sentences we have equality of $e$.
Case $2.1$: $e=i$. Then all elements satisfying $\psi^{j}_{u}(x)$ have $a$-distance $\leq 4\cdot 7^{j}+2\cdot 7^{j}=6\cdot 7^{j}<7^{j+1}$ from $\bar{u}$. In particular, this includes $u$ itself, and then as $u\not\in B(2\cdot 7^{j+1},\bar{u})$ we have
$B(7^{j+1},\bar{u})^{\ast}\models\exists z(\delta_{6\cdot 7^{j}}(\bar{u},u)\land\rho_{2\cdot 7^{j}}(\bar{u},z)\land\psi^{j}_{u}(z)\land\psi^{j}_{\bar{u}}(\bar{u}))$
and so by the forth hypothesis
$B(7^{j+1},\bar{v})^{\ast}\models\exists z(\delta_{6\cdot 7^{j}}(\bar{u},u)\land\rho_{2\cdot 7^{j}}(\bar{u},z)\land\psi^{j}_{u}(z)\land\psi^{j}_{\bar{u}}(\bar{u}))$
(again modulo $g(j+1)$ being sufficiently large). Then (by witnessing) we have $v\in N$ such that:
$2\cdot 7^{j}<d(\bar{v},v)\leq 6\cdot 7^{j}$
$B(7^{j},v)^{\ast},v\equiv^{s}_{g(j)}B(7^{j},v)^{\ast},v)$
$B(7^{j},\bar{u})^{\ast},\bar{u}\equiv^{s}_{g(j)}B(7^{j},\bar{v})^{\ast},\bar{v}$
Moreover, the underlying sets are disjoint within their respective models and so we immediately have:
$B(7^{j},\bar{u}u)^{\ast},\bar{u}u\equiv^{s}_{g(j)}B(7^{j},\bar{v}v)^{\ast},\bar{v}v$
and therefore $\bar{u}u\mapsto\bar{v}v\in I_{j}$.
Case $2.2$: $e<i$. Now $N^{\ast}\models\exists x_{1}...\exists x_{e+1}\gamma_{e+1}$ so there is an element $v\in N$ such that $B(7^{j},\bar{v})\cap B(7^{j},v)=\varnothing$ and $N^{\ast}\models\psi^{j}_{u}(v)$, that is $(B(7^{j},u)^{\ast},u)\equiv^{s}_{g(j)}(B(7^{j},v)^{\ast},v)$ and we can proceed as in the previous case.
Remark 25.
It is worth noting that our definition of locality concepts require the appeal to the standard expansions in the statement of the theorem. In the next section we will consider an alternative approach where locality machinery is defined in directly. We will also see in that in such contexts the appeal to standard expansions is still necessary (example 5).
5. Locality in Semirings
As mentioned earlier, locality in non-classical models have also been investigated by Bizière, Grädel and Naaf (2023) who looked at models defined over ordered semirings. The details of the semiring setting, both in syntax and semantics, make a direct comparison somewhat difficult. Additionally, the authors focus directly on the potential existence of strongly equivalent Gaifman normal forms, i.e. the full Gaifman locality theorem, rather than the question of linking the modelling of local formulas and sentences to strong equivalence. Nevertheless, there are some comparisons we can make between our investigation and theirs which help illuminate the nature of locality in the two settings. First, let us briefly introduce ordered semiring semantics.
{defi}
A commutative semiring is an algebra $(A,+,\cdot,0,1)$ with 01010$\not$=10 ≠ 1 such that $(A,+,0)$ and $(A,\cdot,1)$ are commutative monoids, $\cdot$ distributes over $+$ and 00 is an annihilator for $\cdot$. We say a semiring $K$ is naturally ordered iff the relation $a\leq b$ iff $\exists c:a+c=b$ is a partial order. A particularly important subclass the lattice semirings where the $(A,\leq)$ is a bounded distributive lattice with +=+$\lor$+ = ∨ and $\cdot$=$\land$ = ∧ and the min-max semirings where the natural order is a total order with 00 minimum and $1$ maximum.
Given a relational predicate language $\mathcal{P}$ and a naturally ordered semiring $A$ we define a $\mathcal{P}$-structure $M$ as a triple $(M,\langle P^{M}\rangle_{P\in\mathcal{P}},\langle\neg P^{M}\rangle_{P\in\mathcal{P}})$ where:
  • $M$ is a non-empty set,
  • $P^{M}\colon M^{n}\rightarrow A$ for each $P\in\mathcal{P}$
  • $\neg P^{M}\colon M^{N}\rightarrow A$ for each $P\in\mathcal{P}$
We use the term $\mathcal{P}$-literals to refer to both the usual and negated relation symbols applied to some $\bar{m}\in M$. We say a $\mathcal{P}$-structure is model-defining iff for any pair of complementary literals $P,\neg P$ precisely one of the values $P^{M}$ and $\neg P^{M}$ is 0. We say that it tracks only positive information iff for all negative literals $\neg L^{M}\in\{0,1\}$. Given a tuple $\bar{m}\in M$, we extend the interpretations of literals for $\bar{m}$ to a truth valuation $||-||^{M}$ for all classical first-order formulas $\varphi(\bar{x})$ written in negation normal form defined by induction:
$||L(\bar{m})||^{M}$ $=L^{M}(\bar{m}),$
$||m=m||^{M}$ $=1$ $||m=n||^{M}$ =$0\text{ for }m\not=n$
$||\psi(\bar{m})\lor\lambda(\bar{m})||^{M}$ $=||\psi(\bar{m})||^{M}+||\lambda(\bar{m})||^{M}$ $||\psi(\bar{m})\land\lambda(\bar{m})||^{M}$ $=||\psi(\bar{m})||^{M}\cdot||\lambda(\bar{m})||^{M}$
$||\exists x\varphi(\bar{m},x)||^{M}$ $=\sum\limits_{m\in M}||\varphi(\bar{m},m)||^{M}$ $||\forall x\varphi(\bar{m},x)||^{M}$ $=\prod\limits_{m\in M}||\varphi(\bar{m},m)||^{M}.$
Remark 26.
We can point out a number of differences in both syntax and semantics. On the semantic side, in the semiring setting the focus is more on understanding generalised versions of conjunction and disjunction with the existential and universal quantifiers understood as arbitrary indexed versions of these. This contrasts in our setting where we always take the quantifiers as infimum and supremum without in principle requiring a conjunction or disjunction connective. The two understandings naturally align when working with lattice semirings.
On the syntax side, the semiring setting only considers alternative interpretations of familiar classical formulas. By restricting to formula written in negation normal forms the semantics can be set up without a true negation connective, which is instead is applied only at the base atomic level. Formulas are then built on these literals rather than atomics. [^8] We see an echo of these approach in our definition of isomorphism types. Much as negation is only applied to atomics and literals are taken as the union of these formulas we only apply the expanded truth constants to the base formulas $a\mathord{\reflectbox{/}}R(\bar{x})$ and $R(\bar{x})\mathord{\reflectbox{/}}a$, with formulas of the first kind ‘positive’ and the second ‘negative’.
The relationship between positive and negative literals has further restriction placed upon it. Together the model defining restriction and tracking positive information restriction mean that for all $a\in A$
$\neg a=\begin{cases}1&\text{if }a=0\\ 0&\text{if }a>0.\end{cases}$
From the perspective of residuated lattices where we interpret negation $\neg(x):=x\mathord{\reflectbox{/}}0$ this is equivalent to having no-zero divisors, i.e. $\forall a,b\in A$ $a\cdot b=0$ implies $a=0$ or $b=0$.
Pulling this together, the two investigations overlap for models defined over residuated chains with no-zero divisors and where we are only concerned with the negation normal form fragment of the set of formulas.
Even when we restrict attention to the overlap in subject matter, comparing the investigations requires care due to their definition of the concepts needed to discuss locality. The notion of Gaifman graph is essentially the same; two elements in a model are adjacent iff there is some relation symbol $R$ whose interpretation sends some tuple containing the elements to a value strictly greater than 00 [BGN23, Definition 2.4]. This is means distance and local neighbourhood are understood as we did in our positive recovery of Hanf locality and the corollary result by Fagin, Stockmeyer and Vardi - corollary 17. Indeed, Bizière, Grädel and Naaf also prove a version of this locality result [BGN23, Theorem 3.6] which agrees with our own result at the point of overlap and is proved in essentially the same fashion. It is notable that both their result and corollary 17 each apply in their own way to a broader class of models then where the investigations overlap. In the ordered semiring case it is recovered for all fully idempotent semirings and in ours for all well-connected residuated lattices.
When it comes to Gaifman locality there is a sharper departure. This is driven by the more ambitious focus of Bizière, Grädel and Naaf’s investigation where they look to investigate the status of the full Gaifman theorem - that is whether every formula in their language is strongly equivalent to some Gaifman formula built only from local formula and basic local sentences. One consequence of this is that they require that the associated formula $\varphi^{r,\bar{x}}$ takes on the exact value of the original formula rather than simply matching in terms of the $\models$ relation [BGN23, Appendix A]. This leads them to directly introduce new ball quantifiers of the form $Qy\in B^{\mathcal{P}}_{r}(m)$ with the semantics
$||\exists y\in B^{\mathcal{P}}_{r}(m)\varphi(m,y)||^{M}$ $:=\sum\limits_{n\in B(r,m)}||\varphi(m,n)||^{M}$
$||\forall y\in B^{\mathcal{P}}_{r}(m)\varphi(m,y)||^{M}$ $:=\prod\limits_{n\in B(r,m)}||\varphi(m,n)||^{M}$
Similarly to our considerations of needing both positive and negative distance formulas they also directly introduce scattered quantifiers which directly quantify over only elements sufficiently far away from the specified element. Notably, the two investigations do nearly converge to the overlap in their semantics with Bizière, Grädel and Naaf restricting to min-max semirings.
[…]
Corollary 27.
Let $\mathcal{P}$ be a finite relational language and let $M,N$ be finite $\mathcal{P}$-models over a min-max semiring $A$. We expand our language with ball and scattered quantifiers and define local formula and basic local sentences accordingly. Suppose that $M$ and $N$ satisfy the same basic local sentences. Then $M\equiv N$.
Proof 5.1.
For each semiring $\mathcal{P}$-model $M$ we define a classical $\mathcal{P}$-model counterpart $M^{c}$ to have the same underlying set and where for each relation symbol $R\in\mathcal{P}$ we set:
$R^{M^{c}}:=\{\bar{m}\in M:R^{M}(\bar{m})>0\}.$
By a straightforward induction one can check for any negation normal formula $\varphi(\bar{x})$, possibly including the additional ball and scattered quantifiers, that
$M\models\varphi(\bar{m})\text{ iff }M^{c}\models\varphi(\bar{m}).$
In particular, $M$ and $N$ satisfying the same basic local sentences implies $M^{c}$ and $N^{c}$ satisfy the same basic local sentences. We note that the semantics of the added quantifiers is equivalent to the standard notion [BGN23, Definition 4.1], therefore by the classical Gaifman lemma $M^{c}$ and $N^{c}$ are logically equivalent and so $M$ and $N$ are as well.
That equivalence in modelling for the negation normal fragments with the ‘flattened’ classical model can also be used to show that our lemma with the strong equivalence conclusions can fail in the min-max semiring setting. {exa} Consider a predicate language with a single monadic predicate $P$ and the min-max semiring defined on the unit interval $[0,1]$. We define the models $M$ and $N$ whose domain is a single element $s$ with
$P^{M}(s)=\frac{3}{4}$ $P^{N}(s)=\frac{1}{2}.$
On the one hand $||\exists xP(x)||^{M}=\frac{3}{4}$ and $||\exists xP(x)||^{N}=\frac{1}{2}$ so the models are not strongly equivalent. However, the classical models $M^{\ast^{c}}$ and $N^{\ast^{c}}$, that is the classical counterparts to the standard expansions in the language $\mathcal{P}^{\ast}$ are identical. Therefore $M^{\ast^{c}}$ and $N^{\ast^{c}}$ certainly satisfy the same basic local sentences and so by the same induction as in our previous corollary $M^{\ast}$ and $N^{\ast}$ satisfy the same basic local sentences.
We can expand our algebraic language to include residuated connectives by considering the standard Gödel chain on the unit interval $[0,1]_{G}$ where we interpret $\cdot$ and $\rightarrow$ by:
$a\cdot b=min\{a,b\}$ $a\rightarrow b=\begin{cases}1&\text{if }a\leq b,\\ b&\text{otherwise}.\end{cases}$
Defining the models $M$ and $N$ as above, we can show that the appeal to the behaviour of standard expansions is necessary in the Gaifman lemma, even with a coherent notion of local formulas and basic local sentences provided by the directly defined ball and scattered quantifiers. Just as before the two models $M$ and $N$ are not strongly equivalent. Now, in [DGCN18, Lemma 11, Example 12] we find an example of two models similarly defined to ours which are shown to be logically equivalent. Following the same reasoning (with some straightforward adjustments) one can check that our models are logically equivalent even for the semantics with the added ball and scattered quantifiers. In particular, they must satisfy the same basic local sentences.
[…]
6. Queries
After our proof of the initial Hanf’s theorem we remarked that our ability to immediately translate the classical proof to the many-valued setting indicated that our interest should be in these theorems antecedent. Our Gaifman result, where we generalised not the usual Gaifman’s theorem but its main lemma, somewhat reinforces this; in both cases what we obtained is a sufficiency condition for models being strongly elementary equivalent. Our interest then turns to what it means for model pairs to satisfy the condition and how that might look different in a many-valued context. Accordingly, we turn to a topic that both utilises these theorems and demonstrates a few standard model comparisons that do meet these conditions - model queries.
{defi}
[Model Query] An $n$-ary query for $n\geq 0$ on $\mathcal{P}$-models is a mapping $Q$ that associates a $\mathcal{P}$-model $(A,M)$ with a subset of $M^{n}$ closed under isomorphism, i.e. $(A,M)\cong(B,N)$ via $(f,g)\colon(A,M)\rightarrow(B,N)$ then $Q(N)=f[Q(M)]$.
We say that $Q$ is definable (in a logic of type $\mathcal{L}$) iff there is an $\mathcal{P}$-formula $\varphi(\bar{x})$ such that for every $(A,M)$:
$Q(A,M)=\{\bar{m}\in M^{n}:(A,M)\models\varphi(\bar{m})\}.$
Remark 28.
If we fix our algebra $A$ we can consider queries definable in the language $\mathcal{P}^{\ast}$ defined relative to standard models. That is queries $Q$ for which there is a $\mathcal{P}^{\ast}$ formula $\varphi(\bar{x})$ such that for every $\mathcal{P}$-model $M$:
$Q(M)=\{\bar{m}\in M^{n}:M^{\ast}\models\varphi(\bar{m})\}.$
Moreover, for each $a\in A$ we can consider the notion of $a$-definable queries where $Q$ is $a$-definable iff there is an $\mathcal{P}$-formula $\varphi(\bar{x})$ such that for every $\mathcal{P}$-model $M$:
$Q(M):=\{\bar{m}\in M^{n}:||\varphi(\bar{m})||^{M}\geq a\}.$
In the context of standard models these definitions are equivalent in the sense that a query is definable iff for all $a\in A$ it is $a$-definable iff it is $a$-definable for some $a\in A$. This follows from the syntactic translations $\varphi\cdot a$ and $a\mathord{\reflectbox{/}}\varphi$ respectively.
Given our primary interest in models defined over a fixed algebra we will proceed by looking at queries definable relative to standard models. When $n=0$ we use $1$ and 00 in place of $M$ and $\varnothing$ for the possible outputs of a query.
We can formulate two locality criteria for queries that align with our two locality theorems. Recall the relation $M\leftrightarrows_{r}N$ stating that for each isomorphism type $\tau$ of $\mathcal{P}$ structures the number of elements of $M$ and $N$ with $r$-sphere type $\tau$ is the same, defined relative to any (strict) threshold distance metric. {defi} Let $d(x,y)$ be a distance metric for $\mathcal{P}$-models. An $n$-ary query $Q$ on $\mathcal{P}$-models is $d$-Hanf local iff there is a number $r\geq 0$ such that for every pair of $\mathcal{P}$-models $M,N$ $\bar{m}\in M^{n}$ and $\bar{n}\in N^{n}$:
$(M,\bar{m})\leftrightarrows_{r}(N,\bar{n})\text{ implies }\bar{m}\in Q(M)\text{ iff }\bar{n}\in Q(N).$
The smallest $r$ for which the above condition holds is called the Hanf locality rank of $Q$ and denoted by $hr(Q)$.
Hanf locality tends to be most useful for 00-ary queries. For $n>0$ we use Gaifman locality, again defined relative to a distance metric. {defi} Let $d(x,y)$ be a distance metric for $\mathcal{P}$-models. An $n$-ary query $Q$, $n>0$ on $\mathcal{P}$-models is called $d$-Gaifman local iff there exists a number $r\geq 0$ such that for every $\mathcal{P}$-model $M$ and $\bar{m_{1}},\bar{m_{2}}\in M^{n}$:
$B(r,\bar{m_{1}})\cong B(r,\bar{a_{2}})\Rightarrow\bar{a_{1}}\in Q(M)\text{ iff}\bar{a_{2}}\in Q(M).$
Note a key difference between these two is that Hanf local queries are concerned with different structures whilst Gaifman local ones looking within a single structure.
Our two forms of locality each prove a method for proving the inexpressibility of a query. We can establish a corollary to each locality theorem that definable queries Hanf/Gaifman local respectively. Thus, to show a query is not definable it is sufficient to check it is not Hanf/Gaifman local. Of course this is only applicable for the settings we recovered the respective locality result.
Corollary 29.
Let $Q$ be an $n$-ary query on $\mathcal{P}$-models defined over a residuated chain $A$ with co-atom, and $d(x,y)$ a (strict)-threshold distance metric. Every definable query is $d$-Gaifman local.
Proof 6.1.
Let $Q$ be defined by the $\mathcal{P}^{\ast}$-formula $\varphi(\bar{x})$. Let $k=qr(\varphi)$ and $d=g(k)$ where $g$ is the function defined in the proof of the residuated chain Gaifman lemma. Suppose $B(d,\bar{m})\cong B(d,\bar{n})$, we note this extends to an isomorphism $B(d,\bar{m})^{\ast}\cong B(d,\bar{n})^{\ast}$. $M^{\ast}$ trivially satisfies the same basic local sentences as itself. Letting $\varphi^{k,\bar{x}}(\bar{x})$ be a local $\mathcal{P}^{\ast}$-formula. Then $M^{\ast}\models\varphi^{k,\bar{x}}(\bar{m})$ iff $B(k,\bar{m})^{\ast}\models\varphi(\bar{m})$ iff $B(k,\bar{n})^{\ast}\models\varphi(\bar{n})$ iff $M\models\varphi^{k,\bar{x}}(\bar{n})$. Therefore by the Many-valued Gaifman’s lemma $(M,\bar{m})\equiv^{k}_{s}(M,\bar{n})$ and:
$\bar{m}\in Q(M)\text{ iff }M^{\ast}\models\varphi(\bar{m})\text{ iff }M^{\ast}\models\varphi(\bar{n})\text{ iff }\bar{n}\in Q(M).$
Corollary 30.
Let $Q$ be an $n$-ary query on $\mathcal{P}$-models defined over a well-connected bounded residuated lattice $A$. Let $d_{>\bot}$ be the strict $\bot$ threshold metric. Every definable query is $d_{>\bot}$-Hanf local
Proof 6.2.
Let $Q$ be a definable query by $\varphi(\bar{x})$. Letting $k=qr(\varphi)$ we consider $d=3^{k}$ and suppose $(M,\bar{m})\leftrightarrows_{d}(N,\bar{n})$. From corollary 17 we have $(M,\bar{m})\equiv^{k}_{s}(N,\bar{n})$ and so $\bar{m}\in Q(M)$ iff $M\models\varphi(\bar{m})$ iff $N\models\varphi(\bar{n})$ iff $\bar{n}\in Q(N)$.
Remark 31.
We can easily give a definition of local queries that utilises the full power of Hanf locality. We define a query to be threshold local iff $\exists r\geq 0$ such that for all $\mathcal{P}$-models $M,N$ and $\bar{m}\in M$, $\bar{b}\in N$ if $M,\bar{m}$ and $N,\bar{n}$ have for each $d$-sphere type $\iota$ either:
  • Both $(M,\bar{m})$ $(N,\bar{n})$ have the same number of points of $r$-sphere type $\iota$ or both have at least $m$ points that $r$-realise $\iota$.
  • then $\bar{m}\in Q(M)$ iff $\bar{n}\in Q(N)$.
6.1. Examples of locality
As a small flavour of the method in example, we present a many-valued variant of the two standard queries for which this technique is demonstrated in the classical case [GKL+07, Chapter 2]. We work in the theory of undirected weighted graphs, that is our signature is a single binary relation symbol $E$
{exa}
We say that a weighted graph is $a$-connected iff for any pair of vertices $x,y,$ there is a path in the graph from $x$ to $y$ where every edge has weight $>a$. The 00-connectivity query $Q$ defined by:
$Q(M)=\begin{cases}1&\text{ if M is $0$-connected}\\ 0&\text{ o.w.}\end{cases}$
is not first-order definable on finite weighted graphs.
We consider $M$ to be a cycle of length $2m$ where each edge has weight $>0$ and all the remaining edges have weight 00. We let $N$ be two $m$ cycles where each edge has weight $a$ and all remaining edges weight $\bot$ where $m>2d+1$ for arbitrary $d\in\omega$.
Because each cycle has length $>2d+1$ for any element of $M\cup N$ $B(d,x)$ is always a chain of length $12d+$ 1 where each edge has weight $a$ and all others are weight $\bot$. Therefore every element of $M\cup N$ realises the same $a,d$-sphere type and so $M\leftrightarrows_{d}N$ but $Q$ does not agree on $M$ and $N$ so $Q$ is not Hanf local.
{exa}
The $a$-transitive closure query is not first order definable, where this query is defined by
$Q(M)=\{(m,n)\in M^{2}:n\text{ is in the }a\text{-transitive closure of }m\}.$
Let $d\in\omega$ and consider $M$ to be a directed $a$-chain of length $24d+$ 2 and $m$ as the element in position $12d+$ 1, and $n$ the element in the $33d+$ 3 position. Then $B_{a}(d,(m,n))\cong B_{a}(d,(n,m))$ as each is just the disjoint union of $2$ directed $a$-chains of length $12d+$ 1 centred on $m/n$ respectively, but $(m,n)\in Q(M)$ and $(n,m)\not\in Q(M)$ so $Q$ is not Gaifman local.
7. Conclusions and Further Study
[…]
There remain many open questions. Perhaps the most pressing would be to better understand to what extent the restrictions we required to recover the theorems are necessary and to seek potential counterexamples to the Gaifman Lemma for models defined over residuated lattices that are not well-connected or unwitnessed models generally. During our comparison to locality for semiring semantics we highlighted the question of whether the full Gaifman theorem might be recovered by following the proof strategy based on quantifier elimination utilised there. The presence of residuated connectives poses a serious barrier to adapting the proofs methods used in [BGN23], as unlike the min-max operations they are not monotone in both arguments. Accordingly, a more modest possibility would be to consider a syntax restriction and focus only on so-called ‘positive’ sentences defined with similar inspiration as they are in classical model theory.
Notably, in this investigation we focused solely on finite models. Whilst a natural setting for a locality investigation given the relative importance of locality results in finite model theory [EF95, Lib04], locality concepts can certainly be sensible applied to the arbitrary potentially infinite case. Indeed, both the classical Hanf [Han65] and Gaifman [Gai82] results were proved for arbitrary models, the natural question then is whether a similar investigation could be carried out for arbitrary models and how critical our restriction to finite models in the preceding work. An immediate cause for concern is that for arbitrary models the $k$-isomorphism types are no longer well defined, requiring an infinite conjunction/disjunction. Classically, this is no issue as the set of formulas in a bounded number of variables is finite up to logical equivalence is finite, but an equivalent result has not been established in the generalised setting. This question is additionally interesting as it links to understanding the situations in which the reverse direction of the Ehrenfeucht-Fraïssé theorem can be recovered, something not guaranteed in the residuated lattice setting [DGCN18, Example 23]. Alternatively, one could look to quantifier elimination methods as used by both Gaifman in his original classical proof and Bizière, Grädel and Naaf for semirings.
Acknowledgment
We were partially supported by the Australian Research Council grant DE220100544.